### v1

# references.bib

Abstract.
In this paper, we develop a systematic framework to study the dispersion surfaces of Schrödinger operators $-\Delta+V$, where the potential $V\in C^{\infty}(\mathbb{R}^{n},\mathbb{R})$ is periodic with respect to a lattice $\Lambda\subset\mathbb{R}^{n}$ and respects the symmetries of $\Lambda$. Our analysis relies on an abstract result, that expands on a seminal work of Fefferman–Weinstein [feffer12]: if an operator depends analytically on a parameter, then so do its eigenvalues and eigenprojectors away from a discrete set. As an application, we describe the generic structure of some singularities in the band spectrum of Schrödinger operators invariant under the three-dimensional simple, body-centered and face-centered cubic lattices.
1. Introduction
Analyzing the behavior of waves in periodic structures is a central theme in condensed matter physics, electromagnetism and photonics. This includes for instance electronic conduction: the flow of electrons through a crystal. In the framework of quantum mechanics, these waves solve the time-dependent Schrödinger equation
(1.1) $i{\partial}_{t}\psi=(-\Delta+V)\psi,\qquad\text{where:}$:
  • the potential $V$ is periodic with respect to a lattice $\Lambda$;
  • the function $\psi$ is the wavefunction of the electron, i.e. $|\psi(t,x)|^{2}$ is the density of probability of finding the electron at position $x$, at time $t$.
Solutions of (1.1) can be written as superpositions of time-harmonic waves: functions of the form $e^{-i\mu t}\phi(x)$, where $\phi$ and $\mu$ solve the eigenvalue problem
(1.2) $\mu\phi=(-\Delta+V)\phi.$
The equation (1.2) will be the main focus of this work.
Because the potential $V$ is periodic, the operator $-\Delta+V$ has absolutely continuous spectrum on $L^{2}$, see [reed, Theorem XIII.100]. The corresponding generalized eigenstates are superpositions over $k\in{\mathbb{R}}^{n}$ of Floquet–Bloch modes: solutions $\phi(x;k)$ to
(1.3) $\begin{gathered}(-\Delta+V)\phi(x;k)=\mu(k)\phi(x;k),\quad x\in{\mathbb{R}}^{n},\\ \phi(x+v;k)=e^{ik\cdot v}\phi(x;k),\quad v\in\Lambda.\end{gathered}$
For each $k\in\mathbb{R}^{n}$, the problem (1.3) has a discrete set of solutions $\mu(k)$, which corresponds to the spectrum of $-\Delta+V$ on the space of quasiperiodic functions
(1.4) $L^{2}_{k}=\big{\{}f\in L^{2}_{\mathrm{loc}}(\mathbb{R}^{n},\mathbb{C}):\;f(x+v)=e^{ik\cdot v}f(x),\ v\in\Lambda\big{\}}.$
The maps $k\mapsto\mu(k)$ are called dispersion surfaces; brought together they form the band spectrum of $-\Delta+V$. The local properties of these maps control the effective dynamics of wavepackets [allaire], and singularities in the band spectrum trigger unusual behavior of waves. For instance, Dirac cones – conical intersection of dispersion surfaces in 2D – give rise to Dirac-like propagation of wavepackets: this explains the relativistic behavior of electrons observed in graphene [feffer12, feffer14].
The mathematical analysis of band spectrum singularities started with the seminal work of Fefferman–Weinstein [feffer12], who proved genericity of Dirac points in honeycomb lattices. This has sparked various mathematical investigations of spectral degeneracies in other two-dimensional lattices [finite, lieb, photonic, cassier, chaban]; the only three-dimensional work so far is [guo]. These works share a common strategy, split in two parts:
  • (a)
    proving results for small potentials via perturbation theory and symmetry arguments;
  • (b)
    extending them to generic potentials using an analyticity argument due to [feffer12].
To prove (b), the above works have referred to [feffer12] for details. This motivates the development of a general formalism under which one can apply the Fefferman–Weinstein theory. In particular, [guo] proved that the band spectrum of Schrödinger operator with small potentials, periodic with respect to the body-centered cubic lattice and invariant under the octahedral group, presents a three-fold Weyl point – see Definition 1 and Figure 1. In [guo, §5.2], they conjectured that this extends to large potentials; we prove this conjecture in the present work. In addition, we provide generic results on the band spectrum singularities of Schrödinger operators invariant under the simple, face-centered and body-centered cubic lattices.
1.1. Main results
We formulate here our two main results. The first one describes the behavior of eigenvalues and eigenprojectors of general families of analytic operators. The second one studies band spectrum singularities for Schrödinger operators invariant under cubic lattices. We will need the following two assumptions:
Assumption A1: ${\mathcal{H}}$ is a Hilbert space, ${\mathcal{D}}$ is a dense subspace of ${\mathcal{H}}$, and $T(z),z\in\mathbb{C}$ is a family of closed operators on ${\mathcal{H}}$ with domain ${\mathcal{D}}$ such that:
  1. (i)
    $T(z)$ is analytic in $z$: $\forall f\in{\mathcal{D}}$, $g\in{\mathcal{H}}$, $z\mapsto\langle T(z)f,g\rangle$ is an analytic function on $\mathbb{C}$;
  2. (ii)
    For each $z\in{\mathbb{R}}$, $T(z)$ is self–adjoint and has discrete spectrum;
  3. (iii)
    There exists $C>0$ such that $||T^{\prime}(z)||\leq C$ for all $z\in{\mathbb{C}}$.
Assumption A2: $\mu:(a,b)\rightarrow{\mathbb{R}}$ is a continuous function on an open interval $(a,b)\subset{\mathbb{R}}$, such that for all $z\in(a,b)$, $\mu(z)$ is an eigenvalue of $T(z)$ of constant multiplicity $m$.
Our main result is:
Theorem 1.
Let $T(z)$ satisfy 1.1 and let $\mu:(a,b)\rightarrow{\mathbb{R}}$ satisfy 1.1. There exist a discrete set $D\subset{\mathbb{R}}$ and an open connected neighborhood $U\subset{\mathbb{C}}$ of ${\mathbb{R}}\setminus D$ containing $(a,b)$, such that $\mu$ extends to an analytic function on $U$ and:
  • $\mu(z)$ is an eigenvalue of $T(z)$ of multiplicity $m$ for all $z\in U$;
  • The eigenprojector $\pi(z)$ associated to $\mu(z)$ is analytic in $z\in U$.
Theorem 1 allows us to develop a systematic framework for the generic analysis of dispersion surfaces of periodic Schrödinger operators of the form $H_{z}=-\Delta+zV$. The most interesting case consists of potentials $V$ that are invariant under the lattice symmetries: for instance, $32\pi/$ 3-invariant and even in the case of the honeycomb lattice; or invariant under the octahedral group for cubic lattices. This is because additional symmetries come with higher multiplicities of eigenvalues, which in turn translate to singularities in the band spectrum; this induces exotic behavior of waves. As an application of Theorem 1, we study band spectrum singularities for Schrödinger invariant under the three three-dimensional cubic lattices: the simple, body-centered and face-centered cubic lattices – see §3-4 for precise definitions.
Definition 1.
Let $E\in\mathbb{R},K\in{\mathbb{R}}^{3}$. We say that a Schrödinger operator $H$ has:
  • An $m$-fold quadratic point at $(K,E)$ if $E$ is a $L^{2}_{K}$-eigenvalue of $H$ of multiplicity $m$ and the Floquet–Bloch problem (1.3) has $m$ solutions $\mu_{1}(k),\dots,\mu_{m}(k)$ such that
    (1.5) $\mu_{j}(K+\kappa)=E+{\mathcal{O}}(|\kappa|^{2}),\quad j=1,\dots,m,\quad\kappa\rightarrow 0.$
  • A (two-fold) valley point at $(K,E)$ if $E$ is a double $L^{2}_{K}$-eigenvalue of $H$ and the Floquet–Bloch problem (1.3) has 2 solutions $\mu_{+}(k),\mu_{-}(k)$ such that for some $v\neq 0$:
    (1.6) $\mu_{\pm}(K+\kappa)=E\pm v\cdot\kappa+{\mathcal{O}}(|\kappa|^{2}),\quad\kappa\rightarrow 0.$
  • A three-fold Weyl point at $(K,E)$ if $E$ is a triple $L^{2}_{K}$-eigenvalue of $H$ and the Floquet–Bloch problem (1.3) has 3 solutions $\mu_{1}(k),\mu_{2}(k),\mu_{3}(k)$ such that for some $\alpha\neq 0$:
    (1.7) $\mu_{j}(K+\kappa)=E+\lambda_{\alpha,j}(\kappa)+{\mathcal{O}}(|\kappa|^{2}),\quad j=1,2,3,\quad\kappa\rightarrow 0,$
    where $\lambda_{\alpha,j}(\kappa)$ are the three roots of the polynomial $\lambda^{3}-4|\alpha|^{3}\kappa^{2}\lambda+16\operatorname{Im}(\alpha^{3})\kappa_{1}\kappa_{2}\kappa_{3}$.
Figure 1. Cross sections of a valley point (left) and a three-fold Weyl point (right).
See Figure 1. We show that Schrödinger operators periodic with respect to the cubic lattices and invariant under the octahedral group (the symmetry group of the unit square) feature the above exotic points in their band spectrum.
Theorem 2.
For generic potentials $V\in C^{\infty}(\mathbb{R}^{3},\mathbb{R})$ invariant under the octahedral group and periodic with respect to a cubic lattice $\Lambda\subset\mathbb{R}^{3}$, and generic values of $z\in\mathbb{R}$, the band spectrum of $H_{z}=-\Delta+zV$ has at least:
  1. (i)
    Two three-fold quadratic points if $\Lambda$ is the simple cubic lattice;
  2. (ii)
    One three-fold Weyl point as well as one two-fold and one three-fold quadratic point if $\Lambda$ is the body-centered cubic lattice;
  3. (iii)
    One valley point if $\Lambda$ is the face-centered cubic lattice.
In Theorem 2, “generic potentials $V\in C^{\infty}(\mathbb{R}^{3},\mathbb{R})$” means all of $C^{\infty}(\mathbb{R}^{3},\mathbb{R})$ but a finite union of hyperplanes, and “generic values of $z\in\mathbb{R}$” means all of $\mathbb{R}$ except a closed countable subset of values. In particular, Theorem 2 proves the conjecture formulated in [guo, §5.2].
1.2. Proofs
Our proof of Theorem 1 consists of three main steps:
  • 1.
    Using basic tools from algebraic geometry, we first show that polynomials depending analytically on a parameter admit an analytic factorization when the parameter is away from a discrete set (Theorem 3).
  • 2.
    We then focus on the finite-dimensional analogue of Theorem 1. By applying Step 1 to characteristic polynomials of analytic families of matrices, we perform an analytic diagonalization of such matrices away from a discrete set (Theorem 4).
  • 3.
    Using local Lyapounov–Schmidt procedures, we reduce an analytic family of operators to matrices, to which we can apply Step 2. The challenge is to control the count of eigenvalues within a spectral range to justify the validity of finite-dimensional reductions.
While Theorem 1 is directly inspired by [feffer12], the technical core of its proof relies on different arguments. Both approaches consider maximal extensions of $\mu$ until a critical threshold $b_{*}$ of $z$; they face the same challenge: continue $\mu$ beyond $b_{*}$. Fefferman and Weinstein construct a vector-valued analytic function $F$, specific to the honeycomb setup, whose zeroes characterize where the multiplicity of $\mu$ is no longer $m$. Because the zeroes of $F$ are discrete, this allows them to continue $\mu$ beyond $b_{*}$. We rely instead on a Lyapounov–Schmidt reduction of $-\Delta+b_{*}V$ near the energy $\mu(b_{*})$. This produces a matrix $M(z)$ defined near $b_{*}$, which we analytically diagonalize on a “disk minus one ray” centered at $b_{*}$. This allows us to continue $\mu$ analytically around $b_{*}$.
Theorem 2 follows a well-grounded strategy, rooted again in the work of Fefferman–Weinstein [feffer12]. Given the analytic family of operators $-\Delta+zV$, where $V$ is periodic with respect to a lattice $\Lambda\subset{\mathbb{R}}^{n}$, and $K\in{\mathbb{R}}^{n}$:
  • 1.
    We first describe the multiplicities of $L^{2}_{K}$-eigenvalues of $-\Delta+zV$ for small values of $z$. This combines a general perturbative argument starting from the explicity diagonalizable operator $H_{0}=-\Delta$, with a representation-theoretic argument that relies on the specific symmetries of $V$. Multiplicities higher than one corresponds to intersections of dispersion surfaces at $K$ hence to band spectrum singularities near $K$.
  • 2.
    Because of Theorem 1, the multiplcities of $L^{2}_{K}$-eigenvalues of $-\Delta+zV$ are actually constant for generic values of $z$ – in particular, they coincide with those found in Step 1 for small values of $z$. This extends the results of Step 1 to generic values of $z$.
  • 3.
    The last step studies the situation in the vicinity of $K$, i.e. the $L^{2}_{K+\kappa}$-eigenvalues of $-\Delta+zV$ for $\kappa$ small. This relies again on a perturbative procedure, this time with respect to $\kappa$. We produce an effective equation for these eigenvalues – see e.g. (1.7) in the case of three-fold Weyl points. We derive expressions for the coefficients of the effective equation in terms of the eigenprojector $\pi(z)$ associated to $\mu(z)$ – see Lemma 3.5. This allows us to show that these coefficients are non-zero for generic values of $z$, and to therefore describe qualitatively the band spectrum singularities.
1.3. Relation to existing work
The goal of this paper is to develop a unified framework for the generic analysis of band spectrum singularities in periodic Schrödinger operators. Fefferman–Weinstein [feffer12] showed that honeycomb Schrödinger operators generically have Dirac cones in their band spectrum, i.e. conical singularities; multiple related analyses for other models followed. For instance, in two dimensions, [photonic, cassier] generalized the result of [feffer12] to photonic operators; [lieb] showed that Schrödinger operators invariant under the Lieb lattice have quadratic degeneracies; [chaban] studied the stability of these degeneracies and showed they split to tilted Dirac cones under parity-breaking perturbations. In three dimensions, [guo] showed that Schrödinger operators invariant under the body-centered cubic lattice admit three-fold Weyl points in their band spectrum.
These subsequent works provide a full analysis for small values of $z$ and $\kappa$ (see Steps 1 and 3 in the proof of Theorem 1 outlined in §1.2). They refer to [feffer12] for details about extending their results to generic values of $z$, though their setup is technically different: for instance, [lieb, chaban] identifies quadratic (instead of linear) singularities and [guo] works with triply (instead of doubly) degenerate eigenvalues. Our paper aims to exempt the above works from providing further details, by proving a general statement about the behavior of eigenvalues and eigenprojectors of analytic families of operators: Theorem 1.
1.4. Future projects
The investigation of Dirac cones in honeycomb structures [feffer12] sparked a multitude of mathematical works beside band spectrum singularities: behavior of wavepackets [feffer14], tight-binding analysis [FeffermanLeeWeinsteinTB], emergence of edge states [FeffermanLeeWeinstein, DrouotEdgestates, DrouotWeinstein], propagation of edge states in Dirac systems [BalBeckerDrouot1, BalBeckerDrouot2, DrouotDirac, BalDirac, HXZ], computation of topological invariants [DrouotBEChoney, Ammari2] and Dirac cones in other setups [berk, Ammari1, WeiJunshanHai]. This showcases the importance of Dirac cones in mathematical physics.
The three-dimensional analogue of Dirac cones are Weyl points. We believe that they are the only stable type of spectral degeneracies in three dimensions – see [DrouotWeyl] for an analysis on discrete models. But to the best of our knowledge, one has yet to produce a Schrödinger operator with Weyl points. We plan to use the current paper as a stepping stone. Since band spectrum singularities other than Weyl points are believed to be unstable, they should generically split to Weyl points under perturbations. So adding e.g. a parity-breaking term to the Schrödinger operators discussed in Theorem 2 should produce Weyl points. This belief is reinforced by a two-dimensional analysis of Chaban–Weinstein [chaban], who demonstrated that the quadratic degeneracies of Schrodinger operators invariant under square lattices become Dirac cones after adding an odd potential. Constructing Schrödinger operators with Weyl points has the potential to spark a number of mathematical investigations, such as wavepacket analysis [feffer14], study of surface states (the 3D analogues of edge states), and computation of topological invariants [Monaco].
In [FeffermanLeeWeinsteinTB], the authors show that honeycomb Schrödinger operators converge, in an appropriate sense and in the tight binding limit, to the Wallace model. As an application, they obtain that the set of values of $z$ for which an additional band comes to perturb a Dirac cone is finite. This complements the result of [feffer12], where they proved that this set is discrete. It would be enlightening to perform a similar tight-binding analysis in the case of the three cubic lattices mentioned above, with the analogue of the Wallace model given by the graph Laplacian.
1.5. Acknowledgement
We gratefully acknowledge support from the National Science Foundation DMS-2054589.
2. Spectra of analytic families of operators
2.1. Spectra of analytic families of matrices
In order to find interesting dispersion surfaces of the operator $H$ in (3.4), our strategy will be to look at
(2.1) $H_{z}=-\Delta+zV$
on $L^{2}_{k}$ for varying $k$ and generic values of $z\in{\mathbb{R}}$. For $z$ close to 00, we can understand the spectrum of $H_{z}$ using perturbation theory. We rely on analyticity to analyze $H_{z}$ for $z$ far from 00. One of our primary tools is the following theorem:
Theorem 3.
Let $\{A_{z}\}_{z\in U}$ be an analytic family of monic degree $n$ polynomials in ${\mathbb{C}}[x]$, for some $U\subset{\mathbb{C}}$ open. Then there is a discrete set $D$, such that on any simply connected open set $W\subset U\setminus D$, there exist a decomposition
(2.2) $A_{z}(x)=\prod_{j=1}^{r}(x-\lambda_{j}(z))^{m_{j}},$
where $\lambda_{1},\ldots,\lambda_{r}:W\rightarrow{\mathbb{C}}$ are analytic functions such that $\lambda_{1}(z),\ldots,\lambda_{r}(z)$ are distinct roots of $A_{z}$ and their respective multiplicities $m_{1},\ldots,m_{r}$ are constant on $W$.
An essentially immediate consequence of this theorem is the following result, which we shall apply to operators $T(z)$ satisfying 1.1, restricted to a finite-dimensional subspace:
Theorem 4.
Let $M:U\rightarrow M_{n\times n}({\mathbb{C}})$ be analytic, where $U\subset{\mathbb{C}}$ open, $U\cap{\mathbb{R}}$ is non-empty, and $M(z)$ is Hermitian for all $z\in U\cap{\mathbb{R}}$. Then there exists a discrete set $D$ such that on any simply-connected open set $W\subset U\setminus D$, $W\cap{\mathbb{R}}$ non-empty, $M(z)$ can be decomposed as
(2.3) $M(z)=\sum_{j=1}^{r}\lambda_{j}(z)\pi_{j}(z),\qquad\text{where:}$:
  • $\lambda_{1},\ldots,\lambda_{r}:W\rightarrow{\mathbb{C}}$ are analytic functions such that for all $z\in W$, $\lambda_{1}(z),\ldots,\lambda_{r}(z)$ are distinct eigenvalues of $M(z)$,
  • $\pi_{1},\ldots,\pi_{r}:W\rightarrow M_{n\times n}({\mathbb{C}})$ are eigenprojectors corresponding to $\lambda_{1},\ldots,\lambda_{r}$, respectively, analytic in $z$, and whose ranks are independent of $z$.
Proof.
Let $\chi_{z}$ denote the characteristic polynomial of $M(z)$, using the convention that $\chi_{z}(\lambda)=\operatorname{det}(\lambda-M(z))$. Then $\chi_{z}$ is a monic degree $n$ polynomial, and so by Theorem 3 there exists a discrete set $D$ such that on any simply connected open set $W\subset U\setminus D$, there exist analytic functions $\lambda_{1},\ldots,\lambda_{r}:W\rightarrow{\mathbb{C}}$ such that $\lambda_{1}(z),\ldots,\lambda_{r}(z)$ are the distinct eigenvalues of $M(z)$ for all $z\in W$.
For $z\in W$, let $\pi_{j}(z)$ be the projection onto the generalized eigenspace associated to $\lambda_{j}(z)$. We then claim that $\pi_{j}$ depends analytically on $z$. Fix some $j\in\{1,\ldots,r\}$ and $z_{0}\in W$, and let ${\mathcal{C}}$ be a simple, closed, positively-oriented contour such that $\lambda_{j}(z_{0})$ is the only eigenvalue of $M(z_{0})$ enclosed by ${\mathcal{C}}$. Since each of the functions $\lambda_{\ell}$ are continuous and there are only finitely many of them, there exists $\varepsilon>0$ such that $\lambda_{j}(z)$ is the only eigenvalue of $M(z)$ enclosed by ${\mathcal{C}}$ for all $|z-z_{0}|<\varepsilon$. Consequently, from [spectralproj, Theorem 3.3], $\pi_{j}$ is given on ${\mathbb{B}}_{\varepsilon}(z_{0})$ by
[…]
Since $z_{0}\in W$ was arbitrary, this shows that $\pi_{j}$ is analytic on $W$.
Lastly, observe that for $z\in W\cap{\mathbb{R}}$ (which is assumed to be non-empty), we can write $M(z)$ as in (2.3) since $M(z)$ is Hermitian on this set, and $\pi_{j}(z)\pi_{\ell}(z)=\delta_{j\ell}\pi_{j}(z)$ for $j,\ell\in\{1,\ldots,r\}$ on this set as well. By the identity theorem, these relations must then hold on all of $W$. As a result, for all $j$, $\pi_{j}$ is in fact the eigenprojector corresponding to $\lambda_{j}$, and the corresponding eigenspaces are pairwise orthogonal. Since the multiplicities of the eigenvalues $\lambda_{1},\ldots,\lambda_{r}$ are constant on $W$, we conclude that, for all $j$, $\operatorname{rank}\pi_{j}(z)$ is constant on $W$ as well. ∎
2.2. Proof of Theorem 3
To prove Theorem 3, we use an extension of the notion of a Sylvester matrix.
Definition 2.
Let $P_{j}$ be the vector space of complex polynomials of degree less or equal than $j$, and let $A\in P_{n}$ and $B\in P_{m}$. Define $\phi_{A,B}:P_{m-1}\times P_{n-1}\rightarrow P_{m+n-1}$ by
(2.5) $\phi_{A,B}(Q_{1},Q_{2})=AQ_{1}+BQ_{2}.$
Then $\phi_{A,B}$ is a linear map called the Sylvester matrix associated to $A$ and $B$.
We will generalize the Sylvester matrix as follows. Let $A_{1},\ldots,A_{m}$ be a collection of polynomials for some $m\geq 2$, let $n_{j}=\deg A_{j}$, and let $n=n_{1}+\cdots+n_{m}$. We introduce
(2.6) $\phi_{A_{1},\ldots,A_{m}}$ $:\prod_{j=1}^{m}P_{n-n_{j}-1}\rightarrow P_{n-1}$
(2.7) $(Q_{1},\ldots,Q_{m})\mapsto\sum_{j=1}^{m}A_{j}Q_{j}$
Lemma 2.1.
The image of $\phi_{A_{1},\ldots,A_{m}}$ is the intersection of $P_{n-1}$ with the ideal $(\operatorname{gcd}(A_{1},\ldots,A_{m}))$.
Proof.
It is immediate from the definition of $\phi_{A_{1},\ldots,A_{m}}$ that
(2.8) $\operatorname{im}\phi_{A_{1},\ldots,A_{m}}\subset P_{n-1}\cap(\operatorname{gcd}(A_{1},\ldots,A_{m})).$
To see the reverse containment, by re-indexing the $A_{j}$’s if necessary, we may assume that $\deg(A_{1})\leq\deg(A_{j})$ for $j=2,\ldots,m$. As a result, if $\sum_{j=1}^{m}A_{j}Q_{j}$ has degree less than or equal to $n-1$, then we can divide each of $Q_{2},\ldots,Q_{m}$ by $A_{1}$, so that there exist polynomials $\widetilde{Q}_{j},B_{j}$ such that
(2.9) $Q_{j}=A_{1}B_{j}+\widetilde{Q}_{j}\quad\text{ and }\quad\deg(\widetilde{Q}_{j})<\deg(A_{1})=n_{1}.$
By our assumptions that $m\geq 2$ and $n_{1}=\deg(A_{1})\leq\deg(A_{j})=n_{j}$ for $j\not=1$, we get that $n_{1}\leq\sum_{\ell\not=j}n_{\ell}=n-n_{j}$ for all $j\not=1$. Putting this together, we get that $\deg(\widetilde{Q}_{j})<n-n_{j}$, and so $\widetilde{Q}_{j}\in P_{n-n_{j}-1}$. In addition, we can now write
(2.10) $\sum_{j=1}^{m}A_{j}Q_{j}=A_{1}Q_{1}+\sum_{j=2}^{m}A_{i}(A_{1}B_{j}+\widetilde{Q}_{j})=A_{1}\left(Q_{1}+\sum_{j=2}^{m}A_{j}B_{j}\right)+\sum_{j=2}^{m}A_{j}\widetilde{Q}_{j}.$
From here, observe that since $\widetilde{Q}_{j}\in P_{n-n_{j}-1}$ for $j=2,\ldots,m$, the sum $\sum_{j=2}^{m}A_{j}\widetilde{Q}_{j}$ is a polynomial of degree less than or equal to $n-1$. Consequently, we must have that
(2.11) $\deg\left(A_{1}\left(Q_{1}+\sum_{j=2}^{m}A_{j}B_{j}\right)\right)\leq n-1$
as well, for otherwise the original sum $\sum_{j=1}^{m}A_{j}Q_{j}$ would not be a polynomial of degree less than or equal to $n-1$. Therefore, if we let $\widetilde{Q}_{1}=Q_{1}+\sum_{j=2}^{m}A_{j}B_{j}$, then $\widetilde{Q}_{1}\in P_{n-n_{1}-1}$. Thus we get that
(2.12) $\sum_{j=1}^{m}A_{j}Q_{j}=\sum_{j=1}^{m}A_{j}\widetilde{Q}_{j}\in\operatorname{im}\phi_{A_{1},\ldots,A_{m}},$
and so $\operatorname{im}\phi_{A_{1},\ldots,A_{m}}=P_{n-1}\cap(\operatorname{gcd}(A_{1},\ldots,A_{m}))$, as claimed. ∎
The reason Lemma 2.1 is important is because if we let $B=\operatorname{gcd}(A_{1},\ldots,A_{m})$, then
(2.13) $\operatorname{rank}\phi_{A_{1},\ldots,A_{m}}$ $=\operatorname{dim}((B)\cap P_{n-1})$
(2.14) $=\operatorname{dim}\{Q\in P_{n-1}\;:\;BQ\in P_{n-1}\}$
(2.15) $=\operatorname{dim}\{Q\in P_{n-1}\;:\;\deg Q\leq n-\deg B-1\}$
(2.16) $=n-\deg B.$
The following result is well known:
Lemma 2.2.
Let $M:U\rightarrow M_{n\times n}({\mathbb{C}})$ be analytic, where $U\subset{\mathbb{C}}$ is open and connected. The rank of $M(z)$ is constant except possibly at a discrete set of points in $U$.
The proof relies on identifying a minor determinant which is not uniformly zero of maximal size. As a polynomial in the coefficients of $M$, this function will depend analytically on $z$ and hence its zeroes, which occur where the rank of $M$ decreases by maximality, are isolated.
Using this lemma and our generalization of the Sylvester matrix outlined above, we can now prove the following proposition.
Proposition 2.1.
Let $\{A_{z}\}_{z\in U}$ be an analytic family of monic degree $n$ polynomials in ${\mathbb{C}}[x]$, for some $U\subset{\mathbb{C}}$ open. Then there exists a discrete set $D$ such for $j=1,\ldots,n$,
(2.17) $\rho_{j}=|\{\lambda\text{ is a root of }A_{z}\text{ of multiplicity }j\}|$
is independent of $z\in U\setminus D$
Proof.
For simplicity of notation, we define the following maps:
(2.18) $\phi_{1}=\phi_{A_{z},A^{\prime}_{z}},\quad\phi_{2}=\phi_{A_{z},A^{\prime}_{z},A^{\prime\prime}_{z}},\quad\ldots\quad\phi_{n}=\phi_{A_{z},\ldots,A_{z}^{(n)}},$
where each of the derivatives in (2.18) is being taken with respect to $x$ (as opposed to $z$). Similarly, we define:
(2.19) $B_{1}=\operatorname{gcd}(A_{z},A^{\prime}_{z}),\quad B_{2}=\operatorname{gcd}(A_{z},A^{\prime}_{z},A^{\prime\prime}_{z}),\quad\ldots\quad B_{n}=\operatorname{gcd}(A_{z},\ldots,A_{z}^{(n)}),$
and lastly we define $d_{j}=\deg B_{j}$ for $j=1,\ldots,n$.
Let $n_{j}=\sum_{\ell=0}^{j}\deg A_{z}^{(\ell)}=\sum_{\ell=0}^{j}n-\ell$; then our arguments preceding this proposition tell us that $\operatorname{rank}\phi_{j}=n_{j}-d_{j}$. In addition, observe that $\phi_{k}$ depends analytically on $z$ for $k=1,\ldots,n$, since the entries of $\phi_{k}$, viewed as a matrix, consist either of the coefficients of $A_{z}$ and its derivatives or are otherwise zero. Consequently, by Lemma 2.2 there exists a discrete set $D_{j}\subset U$ such that $d_{j}$ is constant on $U\setminus D_{j}$. Letting $D=\cup_{j=1}^{n}D_{j}$, $D$ is then also discrete as a finite union of discrete sets, and by construction $d_{j}$ is constant on $U\setminus D$ for $j=1,\ldots,n$.
Now let $z_{0}\in U\setminus D$; our assumption that $A_{z}$ is monic implies that we can write
(2.20) $A_{z_{0}}(x)=\prod_{j=1}^{r}(x-\lambda_{j})^{m_{j}},$
where $\lambda_{1},\ldots,\lambda_{r}$ are the $r$ distinct roots of $A_{z_{0}}$, and $m_{1},\ldots,m_{r}$ are the corresponding multiplicities. It then follows that
(2.21) $B_{1}$ $=\prod_{\ell\;:\;m_{\ell}>1}(x-\lambda_{\ell})^{m_{\ell}-1}$
(2.22) $B_{2}$ $=\prod_{\ell\;:\;m_{\ell}>2}(x-\lambda_{\ell})^{m_{\ell}-2}$
[…]
(2.24) $B_{n-1}$ $=\prod_{\ell\;:\;m_{\ell}=n}(x-\lambda_{\ell})$
(2.25) $B_{n}$ $=1.$
A consequence of these formulas is that
(2.26) |$|\{\lambda_{\ell}\text{ is a root of }A_{z_{0}}\;:\;m_{\ell}>j\}|=\deg B_{j}-\deg B_{j+1}=d_{j}-d_{j+1}.$
As a result,
(2.27) $\rho_{j}$ =|$=|\{\lambda\text{ is a root of }A_{z-0}\text{ of multiplicity }j\}|$
(2.28) =|$=|\{\lambda_{\ell}\text{ is a root of }A_{z_{0}}\;:\;m_{\ell}>{j-1}\}|-|\{\lambda_{\ell}\text{ is a root of }A_{z_{0}}\;:\;m_{\ell}>j\}|$
(2.29) $=(d_{j-1}-d_{j})-(d_{j}-d_{j+1}),$
where $d_{0}$ is understood to be $n$ and $d_{j}=0$ for $j>n$. Since $z_{0}$ was arbitrary, this tells us that $\rho_{j}$ is constant on $U\setminus D$ as a sum of constant functions. Therefore the number of distinct roots of $A_{z}$ of any given multiplicity is constant on $U\setminus D$. ∎
Theorem 5.
(Local version) Let $z_{*}\in U\setminus D$. There exists $\varepsilon>0$ such that the conclusion of Theorem 3 holds on $W={\mathbb{B}}_{\varepsilon}(z_{*})$.
Proof.
Let $\lambda_{1},\ldots,\lambda_{r}$ be the distinct roots of $A_{z_{*}}$, and let
(2.30) $\delta=\frac{1}{2}\operatorname{min}_{j\not=\ell}|\lambda_{j}-\lambda_{\ell}|.$
We then say $m\in\{1,\ldots,n\}$ is $\varepsilon$-good if for every root $\lambda$ of $A_{z_{*}}$ of multiplicity $m$, ${\mathbb{B}}_{\delta}(\lambda)$ contains a root of $A_{z}$ of multiplicity $m$ for all $z\in{\mathbb{B}}_{\varepsilon}(z_{*})$.
We claim that for any $m\in\{1,\ldots,n\}$, there exists $\varepsilon>0$ such that $m$ is $\varepsilon$-good. Suppose for contradiction that $m$ is not $\varepsilon$-good for any $\varepsilon>0$; without loss of generality, we may assume that $m$ is maximal with respect to this property. Then for every $\varepsilon>0$, there exists a root $\lambda$ of $A_{z_{*}}$ of multiplicity $m$ and a $z\in{\mathbb{B}}_{\varepsilon}(z_{*})$ such that ${\mathbb{B}}_{\delta}(\lambda)$ does not contain a root of $A_{z}$ of multiplicity $m$. Therefore, if we let $\Sigma$ denote the set of roots of $A_{z_{*}}$ of multipicity $m$, we can construct sequences $(z_{j})_{j\in\mathbb{N}}\subset U\setminus D$, $z_{j}\rightarrow z_{*}$ and $(\lambda_{j})_{j\in\mathbb{N}}\subset\Sigma$ such that for all $j$, $A_{z_{j}}$ does not contain a root of multiplicity $m$ in ${\mathbb{B}}_{\delta}(\lambda_{j})$. Since $A_{z_{*}}$ has only finitely many roots, by passing to subsequences if necessary, we may assume that $\lambda_{j}=\lambda_{*}$ is independent of $j$.
From here, note that $\rho:=|\Sigma|$ is independent of $z$ by Proposition 2.1, and for all $j\in\mathbb{N}$, none of the $\rho$ roots of $A_{z_{j}}$ of multiplicity $m$ are contained in ${\mathbb{B}}_{\delta}(\lambda_{*})$. Therefore, for each $j$, we must either have that there exists a $\tilde{\lambda}_{j}$ such that $A_{z_{j}}$ has two roots in ${\mathbb{B}}_{\delta}(\tilde{\lambda}_{j})$ of multiplicity $m$, or there exists a root $\tilde{\lambda}_{j}$ of $A_{z_{j}}$ of multiplicity $m$ that is not contained in ${\mathbb{B}}_{\delta}(\lambda)$ for any $\lambda\in\Sigma$. Therefore, by again passing to subsequences if necessary, one of two situations must hold:
  • there exists $\tilde{\lambda}_{*}\in\Sigma$ such that for all $j$, $A_{z_{j}}$ has two roots of multiplicity $m$ in ${\mathbb{B}}_{\delta}(\tilde{\lambda}_{*})$,
  • for each $j\in\mathbb{N}$, there exists $\lambda_{j}\not\in\cup_{\lambda\in\Sigma}{\mathbb{B}}_{\delta}(\lambda)$, $\lambda_{j}$ a root of $A_{z_{j}}$ of multiplicity $m$.
Assume the first situation holds. Since the set ${\partial}{\mathbb{B}}_{\delta}(\tilde{\lambda}_{*})$ is compact, $A_{z_{j}}$ converges uniformly to $A_{z_{*}}$ on this set. Because of (2.30), $A_{z_{*}}$ has no zeroes on ${\partial}{\mathbb{B}}_{\delta}(\tilde{\lambda}_{*})$, and therefore uniform convergence tells us that for sufficiently large $j$ and $\zeta\in{\partial}{\mathbb{B}}_{\delta}(\tilde{\lambda}_{*})$,
(2.31) $|A_{z_{j}}(\zeta)-A_{z_{*}}(\zeta)|<\inf_{{\partial}{\mathbb{B}}_{\delta}(\tilde{\lambda}_{*})}|A_{z_{*}}|\leq|A_{z_{*}}(\zeta)|.$
By Rouché’s theorem, $A_{z_{j}}$ and $A_{z_{*}}$ have the same number of zeroes in ${\mathbb{B}}_{\delta}(\tilde{\lambda}_{*})$, counted with multiplicity. However, this is a contradiction, as $A_{z_{j}}$ has at least $2m$ zeroes in this set, counted with multiplicity, while $A_{z_{j}}$ by assumption has precisely $m$ zeroes.
Now assume the second situation holds; we then first show that the sequence $(\lambda_{j})_{j\in\mathbb{N}}$ is bounded. Note that the sequence $(z_{i})$ is bounded by virtue of being convergent, and thus there exists $C>0$ such that $|z_{j}-z_{*}|<C$ for all $j$. In addition, if we write
(2.32) $A_{z}(x)=x^{n}+a_{n-1}(z)x^{n-1}+\cdots+a_{1}(z)x+a_{0}(z),$
then by definition of $A_{z}$ being analytic in $z$, so too are functions $a_{\ell}(z)$ for $\ell=0,\ldots,n-1$. As a result, $a_{\ell}^{\prime}(z)$ exists and is bounded on the closed disk $\overline{{\mathbb{B}}_{C}(z_{*})}$, say by $C_{\ell}$, for $\ell=0,\ldots,n-1$. This implies that for $\ell=0,\ldots,n-1$ and positive integers $j$,
(2.33) $|a_{\ell}(z_{j})-a_{\ell}(z_{*})|\leq C_{\ell}|z_{\ell}-z_{*}|\leq C\cdot C_{\ell}.$
Lagrange’s bound on polynomial roots, together with the reverse triangle inequality, then tells us that for all positive integers $j$,
(2.34) $|\lambda_{j}|\leq\operatorname{max}\left\{1,\sum_{\ell=0}^{n-1}|a_{\ell}(z_{j})|\right\}\leq\operatorname{max}\left\{1,\sum_{\ell=0}^{n-1}|a_{\ell}(z_{*})|+C\cdot C_{\ell}\right\}.$
Consequently the sequence $(\lambda_{j})_{j\in\mathbb{N}}$ is bounded.
Since the sequence is bounded, by again passing to a subsequence if necessary, we may assume that $\lambda_{j}$ converges to some $\lambda_{\infty}$. In addition, since $\lambda_{j}$ is a root of $A_{z_{j}}$ of multiplicity $m$ for all $j$, it is a root of $A_{z_{j}}^{(\ell)}$ for $\ell=0,\ldots,m-1$ (where again the derivative is being taken with respect to $x$, and not $z$). Therefore, for $\ell=0,\ldots,m-1$,
(2.35) $A_{z_{*}}^{(\ell)}(\lambda_{\infty})=A_{\lim_{j\rightarrow\infty}z_{j}}^{(\ell)}\left(\lim_{j\rightarrow\infty}\lambda_{j}\right)=\lim_{j\rightarrow\infty}A_{z_{j}}^{(\ell)}(\lambda_{j})=\lim_{j\rightarrow\infty}0=0.$
Therefore $\lambda_{\infty}$ is a root of $A_{z_{*}}$ of multiplicity at least $m$. However, $d(\lambda_{\infty},\Sigma)\geq\delta$ since $\delta(\lambda_{j},\Sigma)\geq\delta$ by assumption, and therefore by (2.30), $\lambda_{\infty}$ must in fact have multiplicity strictly greater than $m$, which we denote by $m^{\prime}$. Since $m$ is assumed to be the maximal integer that is not $\varepsilon$-good for any $\varepsilon$, it follows that $m^{\prime}$ is $\varepsilon$-good for some $\varepsilon>0$. Therefore, for sufficiently large $j$, ${\mathbb{B}}_{\delta}(\lambda_{\infty})$ contains both $\lambda_{j}$ (which is a root of multiplicity $m$) and a root of multiplicity $m^{\prime}$, distinct from $\lambda_{j}$. However, Rouché’s theorem then implies that $\lambda_{\infty}$ is then a root of $A_{z_{*}}$ of multiplicity $m^{\prime}+m$, which is a contradiction. Consequently the set of $m\in\{1,\ldots,n\}$ which is not $\varepsilon$-good for any $\varepsilon>0$ is empty, thus proving the claim.
Hence there exists a $\varepsilon>0$ such that for all $m\in\{1,\ldots,m\}$, $m$ is $\varepsilon$-good, and by shrinking $\varepsilon$ if necessary, we may assume that ${\mathbb{B}}_{\varepsilon}(z_{*})\subset U\setminus D$. That is, if $z\in{\mathbb{B}}_{\varepsilon}(z_{*})$ and $\lambda_{*}$ is a root of $A_{z_{*}}$ of multiplicity $m$, then ${\mathbb{B}}_{\delta}(\lambda_{*})$ contains precisely one root of $A_{z}$ of multiplicity $m$, which we denote by $\lambda(z)$. Because this holds for every root $\lambda_{*}$ of $A_{z_{*}}$ and the sum of multiplicities of the roots of $A_{z}$ equals $n$, we conclude that $\lambda(z)$ is the only root of $A_{z}$ contained in ${\mathbb{B}}_{\delta}(\lambda_{*})$. Therefore, by the residue theorem,
[…]
Since the integral on the right hand side is analytic in $z$, this implies that $\lambda(z)$ is analytic on ${\mathbb{B}}_{\varepsilon}(z_{*})$. Doing this for each root $\lambda_{*}$ of $A_{z_{*}}$, we get the decomposition (2.2) on ${\mathbb{B}}_{\varepsilon}(z_{*})$, as desired. ∎
We can now prove Theorem 3 using a combination of the local version (Theorem 5) together with the monodromy theorem.
Proof of Theorem 3.
Let $W\subset U\setminus D$ be a simply connected open set and let $z_{0}\in W$. Then by Theorem 5, there exists an open disk $B\subset W$, centered at $z_{0}$ such that the decomposition (2.2) holds on $B$.
Now the monodromy theorem tells us that if there exists an analytic continuation of $\lambda_{j}$ along any path contained in $W$ starting at $z_{0}$, then there exists an extension of $\lambda_{j}$ to all of $W$. In particular, by an analytic continuation of $\lambda_{j}$ along a path $\gamma:[0,1]\rightarrow{\mathbb{C}}$, we mean that, for each $t\in[0,1]$, there exists an open disk $B_{t}$, centered at $\gamma(t)$, and an analytic map $\mu_{t}:B_{t}\rightarrow{\mathbb{C}}$ satisfying
  1. (1)
    $B_{0}=B$ and $\mu_{0}=\lambda_{j}$
  2. (2)
    For each $t\in[0,1]$ there exists $\varepsilon>0$ such that for all $t^{\prime}\in[0,1]$ with $|t-t^{\prime}|<\varepsilon$, $\gamma(t^{\prime})\in B_{t}$ and $\mu_{t^{\prime}}=\mu_{t}$ on $B_{t^{\prime}}\cap B_{t}$.
Therefore, to apply this theorem, let $j\in\{1,\ldots,r\}$, and let $\gamma:[0,1]\rightarrow W$ be a path satisfying $\gamma(0)=z_{0}$. To show that there exists an analytic continuation of $\lambda_{j}$ along the path $\gamma$, we first apply Theorem 5 again to $\gamma(t)$ for each $t\in[0,1]$ to get an open disk $\widetilde{B}_{t}$, centered at $\gamma(t)$, such that the decomposition (2.2) exists on $\widetilde{B}_{t}$. Then for each $t$, define $I_{t}$ to be the connected component of $\gamma^{-1}(\widetilde{B}_{t})$ containing $t$. It follows that $\{I_{t}\}_{t\in[0,1]}$ is an open cover of $[0,1]$, and since this latter set is compact, there exists a finite subcover, which we denote by $\{I_{\ell}\}_{\ell=1}^{s}$ for some positive integer $s$, and for each interval $I_{\ell}$, we denote the corresponding $t$ and open disk $\widetilde{B}_{t}$ by $t_{\ell}$ and $B_{\ell}$, respectively. To ensure that our eventual analytic continuation satisfies property (1) above, we also add the connected component of $\gamma^{-1}(B_{0})$ containing 00, which we denote by $I_{0}$, so that the subcover is now given by $\{I_{\ell}\}_{\ell=0}^{s}$, and $t_{0}=0$ and $B_{0}=B$. Furthermore, by taking a minimal subcover and reindexing and shrinking the sets $I_{\ell}$ for $\ell>0$ if necessary, we may assume that $I_{0}=[0,b_{0})$, $I_{\ell}=(a_{\ell},b_{\ell})$ for $\ell=1,\ldots,s-1$, and $I_{s}=(a_{s},1]$, where $0<a_{1}<\cdots<a_{s}$ and $b_{0}<\cdots<b_{s-1}<1$.
For each of the open disks $B_{\ell}$ we now need analytic maps $\mu_{\ell}:B_{\ell}\rightarrow{\mathbb{C}}$ for $0\leq\ell\leq s$, which we pick inductively. For the base case, simply let $\mu_{0}=\lambda_{j}$. For the inductive case, assume that there exists an analytic map $\mu_{\ell}:B_{\ell}\rightarrow{\mathbb{C}}$ for some $0<\ell<s$ such that $\mu_{\ell}$ is a root of $A_{z}$ for all $z\in B_{\ell}$. Then $B_{\ell}\cap B_{\ell+1}$ must be both nonempty, as it contains $\gamma(I_{\ell}\cap I_{\ell+1})=\gamma((a_{\ell+1},b_{\ell}))$, and connected, as it is the intersection of two disks, and so on $B_{\ell}\cap B_{\ell+1}$ the map $\mu_{\ell}$ must equal precisely one of the roots of the decomposition (2.2) on $B_{\ell+1}$. Thus we can let $\mu_{\ell+1}$ be this root.
To finish the proof of the existence of an analytic continuation of $\lambda_{j}$ along $\gamma$, now let $t\in[0,1]$ be arbitrary. Then $t\in I_{i}$ for some $i$, and consequently there exists an open disk $B_{t}\subset B_{\ell}$ centered at $\gamma(t)$, and we can let $\mu_{t}=\mu_{\ell}|_{B_{t}}$. In addition, since $I_{\ell}$ is open, there exists $\varepsilon>0$ such that $(t\pm\varepsilon)\subset I_{\ell}$. As a result, for all $t^{\prime}\in[0,1]$ satisfying $|t^{\prime}-t|<\varepsilon$, $\gamma(t^{\prime})\in B_{t}\subset B_{\ell}$, from which it follows by construction of the pairs $(B_{\ell},\mu_{\ell})$ for $\ell=0,\ldots,s$, that $\mu_{t^{\prime}}=\mu_{\ell}=\mu_{t}$ on $B_{t^{\prime}}\cap B_{t}$. Thus there exists an analytic continuation of $\lambda_{j}$ along $\gamma$, and since $j\in\{1,\ldots,r\}$ was arbitrary, the monodromy theorem tells us that there exist extensions of $\lambda_{1},\ldots,\lambda_{r}$ to all of $W$. Moreover, these functions must still be roots of $A_{z}$ on $W$, as they are roots along any path in $W$.
To conclude, we note that the functions $z\mapsto\operatorname{mult}(\lambda_{j}(z))$ are locally constant on $W$, and since $W$ is assumed to be simply connected, and thus also connected, it follows that these functions are in fact constant on $W$. In particular, this means that the multiplicities of the roots $\lambda_{1},\ldots,\lambda_{r}$ are constant on $W$. This in turn implies that these roots must be distinct for all $z\in W$, for otherwise their multiplicities would not be constant. ∎
2.3. Analyticity of eigenvalues
In addition to the above tools, the proof our main theorem also requires some techniques from the theory of variation of eigenvalues. In this and the following section, we will restrict our attention to analytic families of closed operators $T(z)$ satisfying 1.1, and a corresponding eigenvalue $\mu:(a,b)\rightarrow{\mathbb{R}}$ satisfying 1.1.
Lemma 2.3.
Let $z_{0},\mu_{0}\in{\mathbb{C}}$ such that $\mu_{0}$ is an eigenvalue of $T(z_{0})$. There exist $\varepsilon,\delta>0$ such that:
  1. (1)
    $\mu_{0}$ is the only eigenvalue of $T(z_{0})$ in ${\mathbb{B}}_{\delta}(\mu_{0})$,
  2. (2)
    for every $z\in{\mathbb{B}}_{\varepsilon}(z_{0})$, $T(z)$ has no eigenvalue on ${\partial}{\mathbb{B}}_{\delta}(\mu_{0})$,
  3. (3)
    for every $z\in{\mathbb{B}}_{\varepsilon}(z_{0})$, the operator
[…]
    is an analytic family of projectors, whose rank is independent of $z$.
The operator $\pi(z)$ defined in (2.37) is typically called the spectral (or Riesz) projector corresponding to ${\partial}{\mathbb{B}}_{\delta}(\mu_{0})$. By evaluating this operator on an eigenvector corresponding to an eigenvalue $\mu$ of $T(z)$ contained in ${\mathbb{B}}_{\delta}(\mu_{0})$ and using Cauchy’s integral formula, one can check that this operator restricts to the identity on the corresponding eigenspace, and in particular the image of this operator contains all eigenspaces corresponding to eigenvalues $\mu$ of $T(z)$ contained in ${\mathbb{B}}_{\delta}(\mu_{0})$.
Proof of Lemma 2.3.
(1)+(2): Since the spectrum of $T(z_{0})$ is discrete, there exists a $\delta>0$ such that $\mu_{0}$ is the only eigenvalue of $T(z_{0})$ contained in $\overline{{\mathbb{B}}_{\delta}(\mu(z_{0}))}$ (so that (1) automatically holds). To prove (2), fix some $\lambda_{0}\in{\partial}{\mathbb{B}}_{\delta}(\mu_{0})$; then, for any $\lambda\in{\partial}{\mathbb{B}}_{\delta}(\mu_{0})$, we can write
(2.38) $T(z)-\lambda$ $=(T(z_{0})-\lambda)\cdot(I+K_{\lambda}(z)),\quad\text{ where }$
[…]
Therefore we seek $\varepsilon>0$ such that for $|z-z_{0}|<\varepsilon$, the operator $K_{\lambda}(z)$ is bounded and has norm less than 1 for all $\lambda\in{\partial}{\mathbb{B}}_{\delta}(\mu_{0})$, from which the claim follows.
Since ${\partial}{\mathbb{B}}_{\delta}(\mu_{0})$ is compact and $(T(z_{0})-\lambda)^{-1}$ is analytic, and thus continuous, in $\lambda$ for all $\lambda\in{\mathcal{C}}$, there exists $M>0$ such that
(2.40) $||(T(z_{0})-\lambda)^{-1}||\leq M$
for all such $\lambda$. Therefore, by the ML inequality,
[…]
for all $\lambda\in{\partial}{\mathbb{B}}_{\delta}(\mu_{0})$. Let
(2.42) $\varepsilon=\frac{1}{CM};$
then $||K_{\lambda}(z)||<1$ for all $|z-z_{0}|<\varepsilon$, and consequently $I+K_{\lambda}(z)$ is invertible. Since $T(z_{0})-\lambda$ is also invertible, we deduce that $T(z)-\lambda$ is invertible as well, thus proving the claim.
(3) This justifies that the operator $\pi(z)$ in (2.37) is well-defined. In addition, it is analytic since its integrand is analytic for all $z\in{\mathbb{B}}_{\varepsilon}(z_{0})$. To see that it is a projector for all such $z$, pick $0<\delta_{0}<\delta$ such that ${\mathbb{B}}_{\delta_{0}}(\mu_{0})$ contains the same eigenvalues of $T(z)$ as ${\mathbb{B}}_{\delta}(\mu_{0})$ (which exists since the resolvent set of $T(z)$ is open); then by the residue theorem, $\pi(z)$ is also equal to the integral in (2.37), but with the ${\partial}{\mathbb{B}}_{\delta}(\mu_{0})$ replaced with ${\partial}{\mathbb{B}}_{\delta_{0}}(\mu_{0})$. Let ${\mathcal{C}}={\partial}{\mathbb{B}}_{\delta}(\mu_{0})$ and let ${\mathcal{C}}_{0}={\partial}{\mathbb{B}}_{\delta_{0}}(\mu_{0})$; then, by the first resolvent identity,
[…]
Lastly, to show that the rank of $\pi(z)$ is independent of $z$, we use a lemma due to Kato [kato, Lemma I.4.10]: if $\pi_{1}$, $\pi_{2}$ are two projectors such that $||\pi_{1}-\pi_{2}||<1$, then $\pi_{1}$ and $\pi_{2}$ have the same (potentially infinite) rank. [^1] In Kato’s book this property is shown for projectors on a finite-dimensional vector space. One can check that the proof applies to infinite-dimensional vector spaces as well. It follows that the set $\{z\in{\mathbb{B}}_{\varepsilon}(z_{0})\;:\;\operatorname{rank}\pi(z)=\pi(z_{0})\}$ is non-empty, open, and closed, and thus equal to ${\mathbb{B}}_{\varepsilon}(z_{0})$. ∎
We will later use the following corollary:
Corollary 2.1.
Let $\mu_{0},z_{0}\in\mathbb{R}$ such that $\mu_{0}$ is eigenvalue of $T(z_{0})$ of multiplicity $1$, let $\phi_{0}$ be a normalized eigenvector of $T(z_{0})$ for the eigenvalue $\mu_{0}$, and let $\varepsilon,\delta$ be the quantities produced by Lemma 2.3. For every $z\in(z_{0}\pm\varepsilon)$, $T(z)$ has a single eigenvalue $\mu(z)$ in $[\mu_{0}\pm\delta]$:
(2.48) $\mu(z)=\mu_{0}+(z-z_{0})\cdot\langle\phi_{0},T^{\prime}(z_{0})\phi_{0}\rangle+O(z-z_{0})^{2}.$
Proof.
For $z\in(z_{0}-\varepsilon,z_{0}+\varepsilon)$, $T(z)$ is a self-adjoint operator with discrete spectrum. We can therefore diagonalize it: there exist eigenvalues $\mu_{k}(z)$ and eigenprojectors $\pi_{k}(z)$ such that:
(2.49) $T(z)=\sum_{j=1}^{\infty}\lambda_{k}(z)\pi_{k}(z).$
In particular, by the residue theorem, the projector $\pi(z)$ from (2.37) is given by:
(2.50) $\pi(z)=\sum_{k:\ |\mu_{k}(z)-\mu_{0}|<\varepsilon}\pi_{k}(z).$
We note that $\pi(z_{0})$ has rank one, because $\mu_{0}$ is the only eigenvalue of $T(z_{0})$ in ${\mathbb{B}}_{\delta}(z_{0})$. Therefore, $\pi(z)$ has rank one as well. We deduce from (2.50) that for every $z\in(z_{0}\pm\varepsilon)$, $T(z)$ has a single eigenvalue $\mu(z)$ in ${\mathbb{B}}_{\delta}(\mu_{0})$, and $\pi(z)$ is its corresponding eigenprojector. Moreover, since $T(z)$ is self-adjoint for $z\in(z_{0}\pm\varepsilon)$, we deduce that in fact $\mu(z)\in(\mu_{0}\pm\delta)$.
Consequently, for all $z\in(z_{0}-\varepsilon,z_{0}+\varepsilon)$, we have the following identities:
(2.51) $(T(z)-\mu(z))\pi(z)=0\quad\text{ and }\quad\pi(z)(T(z)-\mu(z))=0.$
Taking the trace of the first identity in (2.51) gives us the formula $\mu(z)=\operatorname{Tr}(\pi(z)T(z)\pi(z))$, and so $\mu$ is analytic. Therefore, we can take the derivative of the first identity in (2.51), multiple by $\pi(z)$ on the left, and apply the second identity to obtain, after some rearrangement, that
(2.52) $\mu^{\prime}(z)\pi(z)=\pi(z)T^{\prime}(z)\pi(z).$
Taking the trace of (2.52) evaluated at $z_{0}$ and using the formula $\pi(z_{0})=\phi_{0}\otimes\phi_{0}$, we obtain
(2.53) $\mu^{\prime}(z_{0})=\operatorname{Tr}(\pi(z_{0})T^{\prime}(z_{0})\pi(z_{0}))=\langle\phi_{0},T^{\prime}(z_{0})\phi_{0}\rangle.$
We conclude using a Taylor expansion of $\mu$ that:
(2.54) $\mu(z)=\mu_{0}+(z-z_{0})\langle\phi_{0},T^{\prime}(z_{0})\phi_{0}\rangle+O(z-z_{0})^{2}.$
2.4. Proof of Theorem 1
We prove our main result: Theorem 1, which predicts that eigenvalues / eigenprojectors of analytic families of operators are also analytic after taking away a discrete set.
$z_{-1}$$z_{*}$$z_{1}$$z_{2}$${\mathbb{C}}$${\mathbb{R}}$
Figure 2. A possible configuration of the sets $V$ and $D$ (where in this figure $D=\{z_{-1},z_{*},z_{1},z_{2}\}$) whose existences are guaranteed by Theorem 1.
Let ${\mathbb{P}}_{\varepsilon}(z)$ denote the disk ${\mathbb{B}}_{\varepsilon}(z)$ with the ray starting at $z$ and pointing in the direction of the positive imaginary axis removed, i.e.
(2.55) ${\mathbb{P}}_{\varepsilon}(z):={\mathbb{B}}_{\varepsilon}(z)\setminus\{z^{\prime}\;:\;\Re(z^{\prime})=z,\mathfrak{I}(z^{\prime})\geq 0\}.$
To prove the above theorem, we first reformulate it using the following definition:
Definition 3.
A value $b_{*}\in[b,+\infty]$ is good (respectively admissible) if there exist sets
  1. (1)
    $D\subset(b,b_{*})$ finite (respectively of the form $\{z_{j}\}_{j\in\mathbb{N}}$ with $z_{j}$ increasing and converging to $b_{*}$)
  2. (2)
    $U\subset{\mathbb{C}}$ a connected neighborhood of $(a,b_{*})\setminus D$ equal to a union of disks ${\mathbb{B}}_{\varepsilon_{z}}(z)$ centered at points $z\in(a,b_{*})\setminus D$ and sets of the form ${\mathbb{P}}_{\varepsilon_{j}}(z_{j})$ (as in (2.55)) centered at $z_{j}\in D$:
    (2.56) $U=\bigcup_{z\in(a,b_{*})\setminus D}{\mathbb{B}}_{\varepsilon_{z}}(z)\cup\bigcup_{z_{j}\in D}{\mathbb{P}}_{\varepsilon_{j}}(z_{j}),$
such that $\mu$ extends to a continuous function on $U\cup D$, and for all $z\in U$:
  1. (a)
    $\mu(z)$ is analytic and an eigenvalue of $T(z)$ of multiplicity $m$,
  2. (b)
    the associated eigenprojector $\pi(z)$ is analytic.
Good and admissible values in $[-\infty,a]$ are defined analogously.
Using this terminology, we then prove a slightly stronger statement, of which Theorem 1 is a consequence.
Theorem 6.
[…]
We shall prove that +$\infty$+ ∞ is admissible; the proof for $\infty$- ∞ is identical.
Lemma 2.4.
$b$ is good.
Proof.
Let $\pi(z)$ denote the eigenprojector corresponding to $\mu(z)$, and let $z_{0}\in(a,b)$ be arbitrary. Then by Lemma 2.3, there exist $\varepsilon,\delta>0$ such that $\mu(z_{0})$ is the only eigenvalue of $T(z_{0})$ in ${\mathbb{B}}_{\delta}(\mu(z_{0}))$ and for every $z\in{\mathbb{B}}_{\varepsilon}(z_{0})$, $T(z)$ has no eigenvalue on ${\partial}{\mathbb{B}}_{\delta}(\mu(z_{0}))$. By shrinking $\varepsilon$ if necessary, we may also assume $(z_{0}\pm\varepsilon)\subset(a,b)$. Since $\mu$ is assumed to be continuous, we then must have that $\mu((z_{0}\pm\varepsilon))\subset{\mathbb{B}}_{\delta}(\mu(z_{0}))$, for otherwise there would exist $z\in(z_{0}\pm\varepsilon)$ such that $\mu(z)\in{\partial}{\mathbb{B}}_{\delta}(\mu(z_{0}))$, contradicting Lemma 2.3 (2).
Let $\widetilde{\pi}(z)$ be the operator defined in (2.37), so that $\widetilde{\pi}(z)$ is the spectral projector corresponding to ${\partial}{\mathbb{B}}_{\delta}(\mu(z_{0}))$. Since $T(z_{0})$ is self-adjoint, $\widetilde{\pi}(z_{0})$ is precisely the eigenprojector corresponding to $\mu(z_{0})$, and therefore has rank $m$. This in turn implies that $\pi(z)$ has rank $m$ for all $z\in(z_{0}\pm\varepsilon)$ and since $\mu(z)$ has multiplicity $m$ on this set, we deduce that $\widetilde{\pi}(z)$ is in fact the eigenprojector corresponding to $\mu$ on $(z_{0}\pm\varepsilon)$, and thus equal to $\pi(z)$ on this set. Since the integrand of (2.37) is analytic in $z$, this implies $\pi(z)$ is real analytic on $(z_{0}\pm\varepsilon)$, and since $z_{0}\in(a,b)$ was arbitrary, this in fact implies that $\pi(z)$ is real analytic on all of $(a,b)$.
This also justifies that $\mu$ can be locally expressed as
[…]
and since the multiplicity $m$ is assumed to be constant, this tells us that $\mu(z)$ is real analytic on $(a,b)$ as well. Now, since $\mu$ and $\pi$ are both real analytic on $(a,b)$, there exists a connected neighborhood $W\subset{\mathbb{C}}$ of $(a,b)$ such that both $\pi$ and $\mu$ have analytic extensions to $U$. It then will follow that $b$ is good with $D=\emptyset$ if we can find a subset $U\subset W$ of the form (2.56) such that $\mu(z)$ is an eigenvalue of $T(z)$ of multiplicity $m$ and $\pi$ is its associated eigenprojector for all $z\in U$.
Note that, for $z\in(a,b)$,
(2.58) $\pi(z)^{2}-\pi(z)=0\quad\text{ and }\quad(T(z)-\mu(z))\pi(z)=0.$
By the identity theorem, the identities in (2.58) must then be true for all $z\in W$. The first identity tells us that $\pi(z)$ is a projector for all $z\in W$, which, together with the second identity, tells us that $\mu(z)$ is an eigenvalue and the image of $\pi(z)$ is contained in its corresponding eigenspace. Moreover, since $\pi(z)$ is analytic, we can again use [kato, Lemma I.4.10] to conclude that its rank is constant, and therefore equal to $m$ on all of $W$. As a result, $\mu(z)$ is an eigenvalue of multiplicity at least $m$ for all $z\in W$.
To finish, recall that for $z_{0}\in(a,b)$, there exists $\varepsilon>0$ such that $\pi$ is given by (2.37) on ${\mathbb{B}}_{\varepsilon}(z_{0})\subset U$, and as such, its image contains the eigenspace corresponding to $\mu(z)$, and thus the dimension of this eigenspace is bounded above by $m$. Therefore the multiplicity of $\mu(z)$ must be exactly equal to $m$ on ${\mathbb{B}}_{\varepsilon}(z_{0})$. Since such a set exists for all $z_{0}\in(a,b)$, we can then define $U$ to be the union of these disks, and then $\mu$ restricted to $U$ will be the desired extension. Therefore $b$ is good. ∎
Lemma 2.5.
Let $b_{*}\in[b,+\infty]$ be good, and let $D,U$ be the sets associated to $b_{*}$ through Definition 3. Then $\mu$ is $C$-Lipschitz on $U\cup D$.
Proof.
By definition of $b_{*}$ being good, we have that, for all $z\in U$, $\mu$ and $\pi(z)$ are an eigenvalue and corresponding eigenprojector, respectively, of $T(z)$, and as such satisfy the identities in (2.51), and consequently $\mu^{\prime}(z)\pi(z)=\pi(z)T(z)\pi(z)$. Taking the norm of both sides of this equation, it follows from sub-multiplicativity that, for $z\in(a,b)$
(2.59) $|\mu^{\prime}(z)|\leq||T^{\prime}(z)||\leq C.$
Therefore $\mu$ is $C$-Lipschitz on $U$. Since $D$ is finite though, it follows that $U$ is dense in $U\cup D$, and thus we can conclude that $\mu$ is in fact $C$-Lipschitz on $U\cup D$. ∎
Lemma 2.6.
If $b_{0}\in[b,+\infty)$ is good or admissible, then there exists $b_{1}>b_{0}$ good. In particular, all finite admissible values are good.
Proof.
let $D,U$ be the sets associated to $b_{0}$ through Definition 3. Then by Lemma 2.5, $\mu$ is $C$-Lipschitz on $U\cup D$, and therefore extends to a $C$-Lipschitz function (which we shall also denote by $\mu$) on $U\cup D\cup\{b_{0}\}$.
We now apply Lemma 2.3 to $z_{0}=b_{0}$ and $\mu_{0}=\mu(b_{0})$ (which is an eigenvalue of $T(b_{0})$ by continuity of eigenvalues). Let $\widetilde{\pi}(z)$ be the operator defined in (2.37), so that for some $\varepsilon,\delta>0$ and for $z\in{\mathbb{B}}_{\varepsilon_{0}}(b_{0}))$, $\widetilde{\pi}(z)$ is the spectral projector corresponding to eigenvalues contained in ${\mathbb{B}}_{\delta}(\mu(b_{0}))$. Let ${\mathcal{E}}(z)=\operatorname{Im}\widetilde{\pi}(z)$, so that ${\mathcal{E}}(z)$ is a finite-dimensional vector space of dimension independent of $z$. Then, since $T(z)$ is assumed to be acting on a Hilbert space ${\mathcal{H}}$, we can now decompose $T(z)$ with respect to ${\mathcal{E}}(z)\oplus{\mathcal{E}}(z)^{\perp}$:
(2.60) $T(z)=\begin{pmatrix}T_{11}(z)&T_{12}(z)\\ T_{21}(z)&T_{22}(z)\end{pmatrix}.$
Observe that for $z\in(b_{0}\pm\varepsilon)$, $T_{12}(z)=T_{21}(z)=0$ since $T(z)$ is self-adjoint for $z\in{\mathbb{R}}$, and since these operators are analytic, the identity theorem tells us that they must be identically zero on ${\mathbb{B}}_{\varepsilon}(b_{0})$.
In addition, note that by construction, $T_{22}(b_{0})=T(b_{0})|_{{\mathcal{E}}(b_{0})^{\perp}}$ has no eigenvalues in ${\mathbb{B}}_{\delta}(\mu(b_{0}))$, and since $T(b_{0})$ is self-adjoint, this implies $T(b_{0})-\mu(b_{0})$ is invertible, and its norm is bounded by $1/\delta$. By writing
(2.61) $T_{22}(z)-\lambda$ $=(T_{22}(b_{0})-\mu(b_{0}))\cdot(I+K_{\lambda}(z)),\quad\text{ where }$
(2.62) $K_{\lambda}(z)$ $=(T_{22}(b_{0})-\mu(b_{0}))^{-1}(T_{22}(z)-T_{22}(b_{0})+\lambda-\mu(b_{0})),$
we obtain, by a Neumann series argument such as the one following (2.38), that $T_{22}(z)-\lambda$ is invertible for $|z-b_{0}|<\delta/2C$ and $|\lambda-\mu(b_{0})|<\delta/2$.
Let $\varphi_{1},\ldots,\varphi_{n}$ be a basis of ${\mathcal{E}}(b_{0})$. For $|z-b_{0}|<\varepsilon_{0}<\operatorname{min}\{\varepsilon,\delta/2C\}$, $\varepsilon_{0}$ sufficiently small, the set
(2.63) $\{\varphi_{j}(z):=\pi(z)\varphi_{j}\}_{j=1}^{n}$
forms a basis for ${\mathcal{E}}(z)$. Let $M(z)$ be the matrix of $T_{11}(z)$ with respect to this basis. Then $M(z)$ is Hermitian for $z\in(b_{0}\pm\varepsilon_{0})$, and its entries are given by
(2.64) $(M(z))_{ij}=\frac{\langle T(z)\varphi_{i}(z),\varphi_{j}(z)\rangle}{||\varphi_{j}(z)||^{2}},$
from which it follows that $M(z)$ is analytic on ${\mathbb{B}}_{\varepsilon_{0}}(b_{0})$.
By Theorem 4, after possibly shrinking $\varepsilon_{0}$, we may express $M(z)$ for $z\in P:={\mathbb{P}}_{\varepsilon_{0}}(b_{0})$ (as in (2.55)) as
(2.65) $M(z)=\sum_{j=1}^{r}\lambda_{j}(z)\pi_{j}(z),\quad\text{ where:}$:
  • $\lambda_{1},\ldots,\lambda_{r}$ are distinct eigenvalues of $M(z)$ with constant multiplicity that depend analytically on $z$,
  • $\pi_{1},\ldots,\pi_{r}$ are distinct eigenprojectors of $M(z)$ that depend analytically on $z$.
$|f|$${\mathbb{C}}$$U$$U\cap P$$P$$\lambda_{1}(z)$$\lambda_{2}(z)$$\lambda_{3}(z)$$\mu(z)$
Figure 3. A cross section of the dispersion surface $\mu$, and the analytic functions $\lambda_{j}$ describing the eigenvalues of $M(z)$ on $P$. In the situation drawn in this figure, $\mu=\lambda_{2}$ on $U\cap P$.
Note that by Definition 3, and specifically (2.56), $U\cap P$ is connected and contains $(b_{0}-\varepsilon_{0},b_{0})$, and thus is non-empty. Therefore, for $z\in U\cap P$, $\mu(z)$ is an eigenvalue of $T(z)$ in ${\mathbb{B}}_{\delta}(\mu(b_{0}))$, and thus is also an eigenvalue of $M(z)$. As a result, $\mu(z)=\lambda_{j}(z)$ for some $j$ on $U\cap P$, from which it follows that $\mu$ extends to an analytic function on $U\cup P$. This also implies that $\pi$ and $\pi_{j}$ agree on $U\cap P$, and so $\pi$ extends to an analytic function on $U\cup P$ as well. Thus, for all $z\in U\cup W$,
  • $\mu(z)$ is analytic and an eigenvalue of $T(z)$ of multiplicity $m$,
  • $\pi(z)$ is analytic and is the eigenprojector associated to $\mu(z)$.
Let $b_{1}=b_{0}+\varepsilon_{0}$. If $b_{0}$ is good, observe that $U\cup P$ is a connected neighborhood of $(a,b_{1})\setminus(D\cup\{b_{0}\})$ of the form (2.56), therefore $b_{1}$ is good. If on the other hand $b_{0}$ is admissible, then write $D=\{z_{j}\}_{j\in\mathbb{N}}$, with $z_{j}$ increasing to $b_{0}$. Then there exists $J>0$ such that $P$ contains $z_{j}$ for all $j\geq J$. Consequently, in the definition of $U$, we can replace ${\mathbb{P}}_{\varepsilon_{j}}(z_{j})$ with an open disk ${\mathbb{B}}_{\varepsilon_{j}}(z_{j})\subset P$ for some potentially new $\varepsilon_{j}>0$. Then the resulting set $U\cup P$ is a connected neighborhood of $(a,b_{1})\setminus\{z_{1},\ldots,z_{J-1},b_{0}\}$, also of the form (2.56), and so $b_{1}$ is still good. A consequence of this is that $b_{0}$ was in fact good, as can be seen by restricting $\mu$ back to $U$ (after the relevant replacements). Therefore all finite admissible values are good. ∎
Lemma 2.7.
Let $b_{j}\in[b,+\infty)$ be good values, increasing and converging to $b_{*}\in[b,+\infty)$. Then $b_{*}$ is good.
Proof.
By Lemma 2.6, it suffices to prove that $b_{*}$ is admissible. For $j\in\mathbb{N}$, let $D_{j},U_{j}$ be the sets associated to $b_{j}$ through Definition 3, let $\mu_{j}$ denote the corresponding extension of $\mu$, and let $\pi_{j}$ denote the associated eigenprojector. To show that $b_{*}$ is admissible, we then want to construct sets $D,U$ and extensions $\mu_{*}$ of $\mu$ and $\pi_{*}$ of $\pi$ as in Definition 3.
Towards that end, we construct new sets $\widetilde{U}_{j},\widetilde{D}_{j}$ for $j\in\mathbb{N}$ recursively. First, let $\widetilde{U}_{1}=U_{1}$ and $\widetilde{D}_{1}=D_{1}$. Then, for $j\geq 1$, by Definition 3 we can write $U_{j+1}$ as
(2.66) $U_{j+1}=\bigcup_{z\in(a,b_{j+1})\setminus D_{j+1}}{\mathbb{B}}_{\varepsilon_{z}}(z)\cup\bigcup_{z_{\ell}\in D_{j+1}}{\mathbb{P}}_{\varepsilon_{\ell}}(z_{\ell}).$
Let $W_{j+1}$ denote the sub-union consisting of the disks and sets of the form ${\mathbb{P}}_{\varepsilon}(z)$ centered at $z\geq b_{j}$:
(2.67) $W_{j+1}=\bigcup_{z\in[b_{j},b_{j+1})\setminus D_{j+1}}{\mathbb{B}}_{\varepsilon_{z}}(z)\cup\bigcup_{z_{\ell}\in D_{j+1}\cap[b_{j},b_{j+1})}{\mathbb{P}}_{\varepsilon_{\ell}}(z_{\ell}).$
We then define $\widetilde{U}_{j+1}:=\widetilde{U}_{j}\cup W_{j+1}$, from which it follows that $\widetilde{U}_{j+1}$ is of the form (2.56). In addition, $\widetilde{U}_{j}\cap W_{j+1}$ is necessarily connected, and $\mu_{j}$ and $\mu_{j+1}$ agree on this set since $\widetilde{U}_{j}\cap W_{j+1}\subset\widetilde{U}_{j}\cap U_{j+1}$, which is an open and connected set containing $(a,b)$, and therefore $\mu_{j}$ and $\mu_{j+1}$ agree on this latter set by the identity theorem. Thus we get a well-defined extension of $\mu_{j}$ to $\widetilde{U}_{j+1}$, which we now denote by $\mu_{j+1}$. It then follows that $\mu_{j}(z)$ is an eigenvalue of $T(z)$ of multiplicity $m$ for all $z\in\widetilde{U}_{j+1}$ since it satisfies these properties on both $\widetilde{U}_{j}$ and $W_{j+1}$. An identical argument tells us that there exists an analytic extension $\pi_{j+1}$ of $\pi_{j}$ to $\widetilde{U}_{j+1}$ such that $\pi_{j+1}(z)$ is the eigenprojector associated with $\mu_{j+1}(z)$. Lastly, we define $\widetilde{D}_{j+1}=\widetilde{D}_{j}\cup(D_{j+1}\cap[b_{j},b_{j+1}))$; then this set is finite since both $\widetilde{D}_{j}$ and $D_{j+1}$ are finite. Since $\mu_{j}$ is $C$-Lipschitz by Lemma 2.5, $\mu_{j+1}$ extends to a continuous function on $\widetilde{U}_{j+1}\cup\widetilde{D}_{j+1}$.
Now let
(2.68) $U=\bigcup_{j\in\mathbb{N}}\widetilde{U}_{j},\quad D=\bigcup_{j\in\mathbb{N}}\widetilde{U}_{j}.$
Since $\widetilde{U}_{j}\cup\widetilde{D}_{j}\subset\widetilde{U}_{j+1}\cup\widetilde{D}_{j+1}$ for all $j\in\mathbb{N}$, we can then define $\mu_{*}:U\cup D\rightarrow{\mathbb{C}}$ and $\pi_{*}:U\cup D\rightarrow{\mathbb{C}}$ as the direct limit of the extensions $\mu_{j}$ to the sets $\widetilde{U}_{j}\cup\widetilde{D}_{j}$, and $\pi_{j}$ to the sets $\widetilde{U}_{j}$, respectively. As such, it follows that the sets $U$ and $D$ (where $D$ is possibly infinite) together with the extensions $\mu_{*}$ and $\pi_{*}$ satisfy the conditions of Definition 3, and so $b_{*}$ is admissible. However, by Lemma 2.6, this implies that $b_{*}$ is in fact good. ∎
Proof of Theorem 1.
Assume that +$\infty$+ ∞ is not admissible and consider
(2.69) $b_{+}=\operatorname{sup}\{b_{*}\in[b,+\infty)\;:\;b_{*}\text{ admissible}\}.$
Then by Lemma 2.4, this supremum is being taken over a non-empty set, and so it is well-defined. Let $b_{j}$ be an increasing sequence of admissible values converging to $b_{+}$. By Lemma 2.6, $b_{j}$ is good for all $j$, and so by Lemma 2.7, $b_{+}$ is good and therefore also admissible. However, this then implies by Lemma 2.6 that there exists $b_{*}>b_{+}$ that is also good, contradicting maximality of $b_{+}$. Therefore +$\infty$+ ∞ must be admissible. ∎
3. Dispersion Surfaces of Schrödinger Operators: General Theory
The rest of this paper focuses on applications of Theorem 1. Specifically, we develop a framework for analyzing the dispersion surfaces of Schrödinger operators $-\Delta+V$ for generic potentials $V$ invariant under a lattice $\Lambda$, i.e. periodic with respect to $\Lambda$ and symmetric with respect to the point group of $\Lambda$.
This section concentrates on this general set-up. We first review Floquet–Bloch theory and define lattice-invariant potentials. We then state and prove perturbative lemmas on Floquet–Bloch eigenvalues of $-\Delta+zV$. Brought together, these results outline our strategy to describe the generic structure of the dispersion surfaces of invariant Schrödinger operators. In the next section we apply this framework to analyze the dispersion surfaces of Schrödinger operators with potentials invariant under cubic lattices.
3.1. Floquet–Bloch Theory
We begin with a review of lattices and Floquet–Bloch theory. Given a basis $v_{1},\ldots,v_{n}$ of ${\mathbb{R}}^{n}$, the lattice $\Lambda$ generated by $v_{1},\ldots,v_{n}$ is the set $\Lambda={\mathbb{Z}}v_{1}\oplus\cdots\oplus{\mathbb{Z}}v_{n}$. Given $k\in{\mathbb{R}}^{n}$, the space of $k$-quasiperiodic functions with respect to $\Lambda$ is
(3.1) $L^{2}_{k}=\{f\in L^{2}_{\text{loc}}\;:\;f(x+v)=e^{ik\cdot v}f(x)\;\forall v\in\Lambda\}.$
In this context, we refer to $k$ as the quasi-momentum of functions $f\in L^{2}_{k}$. In addition, observe that the space of $\Lambda$-periodic functions is simply $L^{2}_{0}$, and $f\in L^{2}_{k}$ if and only if $e^{-ik\cdot x}f\in L^{2}_{0}$. The correspondence between $L^{2}_{0}$ and $L^{2}_{k}$ then induces an inner product on $L^{2}_{k}$ given by:
[…]
where $\mathbb{R}^{n}/\Lambda$ is a fundamental cell for $\Lambda$. We similarly define Sobolev spaces $H^{s}_{k}$, $s\in\mathbb{N}$ by
(3.3) $H^{s}_{k}=\{f\in L^{2}_{k}\;:\;{\partial}^{\alpha}f\in L^{2}_{k}\;\forall|\alpha|\leq s\}.$
Lastly, we define the dual lattice $\Lambda^{*}$ (also often referred to as the reciprocal lattice) as $\Lambda^{*}={\mathbb{Z}}k_{1}\oplus\cdots{\mathbb{Z}}k_{n}$, where $k_{1},\ldots,k_{n}$ satisfy the relation $k_{i}\cdot v_{j}=2\pi\delta_{ij}$. We then refer to $k_{1},\ldots,k_{n}$ as the dual (or reciprocal) basis.
We consider the Schrödinger operator $H=-\Delta+V$, where $V$ is smooth and periodic with respect to $\Lambda$. The Floquet–Bloch eigenvalue problem at quasi-momentum $k\in{\mathbb{R}}^{n}$ is
(3.4) $\begin{gathered}H\phi(x;k)=\mu(k)\phi(x;k),\quad x\in{\mathbb{R}}^{n},\\ \phi(x+v;k)=e^{ik\cdot v}\phi(x;k),\quad v\in\Lambda.\end{gathered}$
A $L^{2}_{k}$-solution $\phi$ to the above problem is called a Floquet–Bloch state. The operator $H$ is a self-adjoint unbounded operator on $L^{2}_{k}$ (respectively $L^{2}$) with domain $H^{2}_{k}$ (respectively $H^{2}$). Using elliptic regularity, the operator $H$ on $L^{2}_{k}$ has a discrete spectrum; the collection of its eigenvalues, seen as functions of $k$, are called the dispersion surfaces of $H$.
Since the problem (3.4) is invariant under the change $k\mapsto k+k^{\prime}$ for $k^{\prime}\in\Lambda^{*}$, we can restrict our attention to $k$ varying over the Brillouin zone ${\mathcal{B}}$: the set of points $k\in{\mathbb{R}}^{n}$ which are closer to the origin than to any other point of $\Lambda^{*}$. Moreover, we can recover the $L^{2}$-spectrum of $H$ from the $L^{2}_{k}$-spectra for $k\in{\mathcal{B}}$ [reed]:
(3.5) $\sigma_{L^{2}}(H)=\bigcup_{k\in{\mathcal{B}}}\sigma_{L^{2}_{k}}(H).$
3.2. Invariant Potentials
In this section, we fix a lattice $\Lambda$ with basis $v_{1},\ldots,v_{n}$ and reciprocal basis $k_{1},\ldots,k_{n}$. Let $G$ denote the point group of the lattice $\Lambda$, namely the subgroup of its isometry group which keeps the origin fixed. Observe that $G$ is necessarily finite: every element $g\in G$ must necessarily send the basis $v_{1},\ldots,v_{n}$ to another basis of ${\mathbb{R}}^{n}$ consisting of vectors in $\Lambda$, and since $g$ is an isometry, we must have that $||gv_{j}||=||v_{j}||$ for $j=1,\ldots,n$, which implies there are only finitely many lattice vectors to which $g$ can send each basis element.
The group $G$ acts isometrically on scalar-valued functions:
(3.6) $g_{*}f(x):=f(g^{\top}x).$
We will later need an induced action of a subgroup $G_{0}$ of $G$ on $L^{2}_{k}$ for some quasi-momentum $k$. However, in order for this action to be well-defined, we need $G_{0}$ to satisfy an additional criterion.
Definition 4.
We say $g\in G$ is $k$-invariant if
(3.7) $gk\in k+\Lambda^{*}.$
Analogously, we say a subgroup $G_{0}$ of $G$ is $k$-invariant if $g$ is $k$-invariant for all $g\in G_{0}$.
To see that $k$-invariant subgroups give well-defined actions, note that if $G_{0}$ is such a subgroup and $g\in G_{0}$, then by definition there exists $k^{\prime}\in\Lambda^{*}$ such that $gk=k+k^{\prime}$. Then for all $v\in\Lambda$, $g^{\top}v\in\Lambda$ as well by definition of $G$, and as a result
(3.8) $g_{*}\psi(x+v)$ $=\psi(g^{\top}x+g\top v)=e^{ik\cdot g^{\top}v}\psi(g^{\top}x)=e^{igk\cdot v}g_{*}\psi(x)$
(3.9) $=e^{i(k+k^{\prime})\cdot v}g_{*}\psi(x)=e^{ik\cdot v}g_{*}\psi(x).$
In particular, this shows that $k$-invariant group elements map $L^{2}_{k}$ to itself.
We now define potentials invariant with respect to $\Lambda$.
Definition 5.
Let $\Lambda$ be a lattice with point group $G$. We say that $V\in C^{\infty}({\mathbb{R}}^{n},{\mathbb{R}})$ is $\Lambda$-invariant if:
  1. 1)
    $V$ is $\Lambda$-periodic, i.e. $V(x+v)=V(x)$ for all $x\in{\mathbb{R}}^{2}$ and $v\in\Lambda$,
  2. 2)
    $V$ is $G$-invariant, i.e. $g_{*}V=V$ for all $g\in G$.
When the lattice $\Lambda$ is clear from the context, we will omit it and simply refer to $V$ as an invariant potential.
When $V$ is an invariant potential, the fact that $V$ is $\Lambda$-periodic enables us to expand $V$ as a Fourier series with coefficients $\{V_{m}\}_{m\in{\mathbb{Z}}^{n}}$:
(3.10) $V(x)$ $=\sum_{m\in{\mathbb{Z}}^{n}}V_{m}e^{i(m_{1}k_{1}+\cdots+m_{n}k_{n})\cdot x}$
(3.11) $V_{m}$ $=\left\langle e^{i(m_{1}k_{1}+\cdots+m_{n}k_{n})\cdot x},V\right\rangle.$
For simplicity of notation, if $k\in\Lambda^{*}$ so that $k=m_{1}k_{1}+\cdots+m_{n}k_{n}$ for some $m\in{\mathbb{Z}}^{n}$, we shall also denote $V_{m}$ by $V_{k}$. If we then view these coefficients as a function on $\Lambda$, they are invariant under an induced action of $G$:
(3.12) $g_{*}V_{k}=V_{g^{\top}k}=\left\langle e^{ig^{\top}k\cdot x},V\right\rangle=\left\langle e^{ik\cdot x},g_{*}V\right\rangle=\left\langle e^{ik\cdot x},V\right\rangle=V_{k}.$
An example of invariant potentials that has been studied extensively is honeycomb lattice potentials: potentials invariant under $32\pi/$ 3-rotations and parity and periodic with respect to the equilateral lattice. For later reference, we now describe two properties of invariant potentials which naturally extend properties of honeycomb lattice potentials.
First, observe that if $V$ is a $\Lambda$-invariant potential and $O$ is an orthogonal transformation, then $V\circ O^{*}$ is an $O\Lambda$-invariant potential. An immediate consequence of this is that the spectral properties of $H_{z}$ on $L^{2}_{k}$ are the same as those of $H_{z}$ on $L^{2}_{gk}$ for all $g\in G$. Together with the $\Lambda^{*}$-periodicity of the Floquet-eigenvalue problem (3.4), this implies that the dispersion surfaces of $H$ near a quasi-momenta $k\in{\mathbb{R}}^{n}$ are determined locally by those near $gk$. Consequently, it suffices to consider quasi-momenta whose orbits under $G$ are distinct.
Second, every lattice $\Lambda$ is necessarily invariant under the negative of the identity, which implies that $-I\in G$. Therefore, by $G$-invariance, every invariant potential $V$ is necessarily even. Together with the assumption that $V$ is real, this implies that if $(\phi(x;k),\mu)$ is an eigenpair of the Floquet–Bloch problem (3.4) with quasi-momentum $k$, then so too is $(\overline{\phi(-x;k)},\mu)$.
3.3. Decomposing $L^{2}_{K}$ via a $K$-Invariant Subgroup
Fix some $K\in{\mathbb{R}}^{n}$; then $\mu_{0}=||K||^{2}$ is an $L^{2}_{K}$-eigenvalue of $-\Delta$. We define a set $[K]$ as follows:
(3.13) $[K]:=\{k\in K+\Lambda^{*}\;:\;||k||=||K||\}.$
Then by Corollary A.2,
(3.14) $m_{-\Delta}(\mu_{0})=\Big{|}[K]\Big{|}.$
For the rest of this section, we make the following assumption on $K$:
Assumption A3: There exists an abelian subgroup $G_{0}$ of $G$ such that $G_{0}K=[K]$ and $|G_{0}|=|G_{0}K|$.
Although this might appear at first glance to be a restrictive assumption, we will see in §4 that in many applications, such a subgroup exists. The reason this assumption is helpful is that by construction of $[K]$, $G_{0}$ is necessarily $K$-invariant, and thus has a well-defined action on $L^{2}_{K}$. In addition, by our assumption that $V$ is $G$-invariant and the fact that $g_{*}$ is the pushforward by an orthogonal matrix for every $g\in G_{0}$, $H$ commutes with the action of $G_{0}$ on $L^{2}_{K}$. We can therefore reduce the spectral problem for $H$ on $L^{2}_{K}$ to spectral problems on the invariant subspaces of $G_{0}$.
Before we perform this reduction, however, we introduce some notation. Let $g_{1},\ldots,g_{\ell}$ denote a minimal system of generators of $G_{0}$, with respective orders $n_{1},\ldots,n_{\ell}$. Since $G_{0}$ is assumed to be abelian, it follows that $G_{0}\cong\bigoplus_{j=1}^{\ell}{\mathbb{Z}}_{n_{j}}$. In addition, if $g\in G_{0}$ is of order $N$, then $\sigma_{L^{2}_{K}}(g_{*})$, the spectrum of $g_{*}$ viewed as an operator on $L^{2}_{K}$, is contained in the $N$-th roots of unity $U_{N}$ (and in fact, we will see in Lemma 3.1 that $\sigma_{L^{2}_{K}}(g_{*})=U_{N}$). This follows first from the fact that $g_{*}$ has finite order, and consequently has pure point spectrum, and if $\omega$ is an eigenvalue of $g_{*}$, then $g^{N}=e$ implies $\omega^{N}=1$, and so $\omega\in U_{N}$. With this in mind, we define:
(3.15) ${\mathbb{J}}:=\prod_{j=1}^{\ell}\{1,\ldots,n_{j}\}\quad\text{ and }\quad{\mathbb{U}}:=\prod_{j=1}^{\ell}U_{n_{j}},$
so that $G_{0}=\{g^{j}\;:\;j\in{\mathbb{J}}\}$, where we are using the multi-index notation $g^{j}=g_{1}^{j_{1}}\cdots g_{\ell}^{j_{\ell}}$.
Again using the fact that $G_{0}$ is abelian, we can then simultaneously diagonalize the operators $(g_{j})_{*}$, which leads us to the following decomposition of $L^{2}_{K}$:
(3.16) $L^{2}_{K}=\bigoplus_{\omega\in{\mathbb{U}}}L^{2}_{K,\omega},\quad L^{2}_{K,\omega}:=\bigcap_{j=1}^{\ell}\operatorname{ker}_{L^{2}_{K}}((g_{j})_{*}-\omega_{j}).$
It is worth noting that the spaces $L^{2}_{K,\omega}$ for $\omega\in{\mathbb{U}}$ are pairwise orthogonal by virtue of the operators $(g_{j})_{*}$ being unitary.
Lastly, it will also simplify our later computations by introducing a convenient method of enumerating elements of $[K]$. Specifically, for each $j\in{\mathbb{J}}$ we define $m(j)\in{\mathbb{Z}}^{n}$ as the $n$-tuple satisfying
(3.17) $g^{j}K=K+m(j)\cdot(k_{1},\ldots,k_{n});$
then $m(j)$ exists and is unique by 3.3.
3.4. Strategy
Our goal is to describe the structure of dispersion relations of
(3.18) $H_{z}=-\Delta+zV,$
near some quasi-momentum $K\in{\mathbb{R}}^{n}$ for generic values of $z$, where we continue to assume that $K$ together with a subgroup $G_{0}$ of $G$ satisfy 3.3. Our strategy relies on the four key lemmas stated below; their proofs are postponed to Section 4.5. We will start with a result of eigenvalues of $-\Delta$ on $L^{2}_{K}$.
Lemma 3.1.
Let $K\in{\mathbb{R}}^{n}$ and $G_{0}$ of $G$ satisfy 3.3. For each $\omega\in{\mathbb{U}}$, $||K||^{2}$ is an $L^{2}_{K,\omega}$-eigenvalue of $-\Delta$ of multiplicity 1, with corresponding normalized eigenvector given by
(3.19) $\phi_{\omega}(x)=\frac{1}{\sqrt{|G_{0}|}}\sum_{j\in{\mathbb{J}}}\omega^{j}e^{ig^{-j}K\cdot x}.$
Lemma 3.1 together with Corollary 2.1 will then help us construct an eigenvalue $\mu(z)$ of $H_{z}$ satisfying 1.1.
Lemma 3.2.
Let $K\in{\mathbb{R}}^{n}$ and $G_{0}$ of $G$ satisfy 3.3 and let $\omega\in{\mathbb{U}}$. There exist $\varepsilon,\delta>0$ such that for $z\in(-\varepsilon,\varepsilon)$, $H_{z}$ has a unique $L^{2}_{K,\omega}$-eigenvalue in $(||K||^{2}\pm\delta)$, given by
(3.20) $\mu(z)=||K||^{2}+z\cdot\sum_{j\in{\mathbb{J}}}\omega^{j}V_{m(j)}+{\mathcal{O}}(|z|^{2}),$
where $m(j)$ is the multi-integer defined in (3.17).
A straightforward application of our main theorem (Theorem 1) implies that $H_{z}$ has an $L^{2}_{K,\omega}$-eigenvalue $\mu(z)$ splitting from the $L^{2}_{K}$-eigenvalue $||K||^{2}$ of $H_{0}=-\Delta$, which for generic $z\in{\mathbb{R}}$ is simple and whose eigenprojector can be extended to an analytic map away from a discrete set.
When $K$ is a vertex of the Brillouin zone, we will be able to compute the generic multiplicities of the $L^{2}_{K}$-eigenvalues of $H_{z}$ splitting from the eigenvalue $||K||^{2}$ of $-\Delta$ using symmetry arguments. We will then describe the structure of the corresponding dispersion surfaces near $K$ using the following three results.
Lemma 3.3.
Let $\mu(z)$ be an $L^{2}_{K}$-eigenvalue of $H_{z}$ for some $z\in{\mathbb{R}}$, let $\pi(z):L^{2}_{K}\rightarrow L^{2}_{K}$ be the corresponding eigenprojector, and let ${\mathcal{E}}(z)$ be the corresponding eigenspace. There exists $\varepsilon,\delta>0$ such that for $|\kappa|<\varepsilon$, the $L^{2}_{K+\kappa}$ eigenvalues of $H_{z}$ in ${\mathbb{B}}_{\delta}(\mu_{0})$ are given by
(3.21) $\mu(z)+\lambda_{j}(z,\kappa)+{\mathcal{O}}(|\kappa|^{2}),$
where $\lambda_{j}(\kappa)$ are the ${\mathcal{E}}$-eigenvalues of $M(z,\kappa):=-\pi(z)(2i\kappa\cdot\nabla)\pi(z)$.
For the next lemma, we continue to let $M(z,\kappa):=-\pi(z)(2i\kappa\cdot\nabla)\pi(z)$, although we no longer assume that $z\in{\mathbb{R}}$, so that we can use this lemma in conjunction with Theorem 1.
Lemma 3.4.
Let $\mu(z)$ be an $L^{2}_{K}$-eigenvalue of $H_{z}$ for some $z\in{\mathbb{R}}$ such that the corresponding eigenprojector $\pi(z)$ depends analytically on $z\in U$. The characteristic polynomial of $M(z,\kappa)$, acting on the finite-dimensional eigenspace ${\mathcal{E}}(z)$, depends analytically on $z\in U$.
Lastly, in order to compute the characteristic polynomial of $M(z,\kappa)$, we will express this matrix with respect to a basis consisting of one vector from each of the subspaces $L^{2}_{K,\omega}$ for $\omega$ in some subset of ${\mathbb{U}}$. This final lemma will help us simplify these computations.
Lemma 3.5.
Let $\phi\in L^{2}_{K,\omega}$ and let $\psi\in L^{2}_{K,\widetilde{\omega}}$ for $\omega,\widetilde{\omega}\in{\mathbb{U}}$. Then for all $j\in{\mathbb{J}}$, $\langle\phi,\nabla\psi\rangle$ is an eigenvector of $g^{j}$ with corresponding eigenvalue $\omega^{-j}\widetilde{\omega}^{j}$. Moreover, if $K$ is a vertex of the Brillouin zone ${\mathcal{B}}$, then $\langle\phi,\nabla\phi\rangle=0$.
3.5. Proofs of Lemmas 3.13.5
Proof of Lemma 3.1.
We first note that $||K||^{2}$, as an $L^{2}_{K}$-eigenvalue of $-\Delta$, by 3.3 has multiplicity $|G_{0}K|=|G_{0}|$. Consequently, it suffices to prove that for each $\omega\in{\mathbb{U}}$, the function $\phi_{\omega}$ in (3.19) is a normalized $L^{2}_{K,\omega}$-eigenvector for the eigenvalue $||K||^{2}$, for then this eigenvalue on $L^{2}_{K,\omega}$ would necessarily be simple since $|{\mathbb{U}}|=|G_{0}|$.
To see that $\phi_{\omega}$ is normalized, observe that the $|G_{0}|$ functions $e^{ig^{j}K\cdot x}$ form an orthonormal system because of 3.3. Indeed, each of these exponentials is distinct, for if $g^{j}K=g^{j^{\prime}}K$, then $g^{j-j^{\prime}}K=K$, which implies $g^{j-j^{\prime}}=I$ because $|G_{0}K|=|G_{0}|$. Therefore $g^{j}=g^{j^{\prime}}$ and $j=j^{\prime}$.
To show that $\phi_{\omega}\in L^{2}_{K,\omega}$, we first note that $\phi_{\omega}\in L^{2}_{K}$ since $g^{j}K\in K+\Lambda^{*}$ for every $j\in{\mathbb{J}}$. In addition, we compute that
(3.22) $(g_{1})_{*}\phi_{\omega}(x)=\frac{1}{\sqrt{|G_{0}|}}\sum_{j\in{\mathbb{J}}}\omega^{j}e^{ig^{-j}K\cdot g^{\top}x}=\frac{\omega_{1}}{\sqrt{|G_{0}|}}\sum_{j\in{\mathbb{J}}}\omega^{j-e_{1}}e^{ig^{-j+e_{1}}K\cdot x}=\omega_{1}\phi_{\omega}(x),$
with similar identities when testing the pushforward operators by $g_{2},\ldots,g_{\ell}$. We conclude by noting that $\phi_{\omega}$ is an eigenvector of $-\Delta$ with eigenvalue $||K||^{2}$ since $||g^{j}K||=||K||$ by virtue of $g^{j}$ being an orthogonal matrix for every $j\in{\mathbb{J}}$. ∎
Proof of Lemma 3.2.
By Lemma 3.1 and Corollary 2.1, there exist $\varepsilon,\delta>0$ such that $H_{z}$ has a single eigenvalue $\mu(z)$ in $(||K||^{2}\pm\delta)$ for all $z\in(-\varepsilon,\varepsilon)$, given by:
(3.23) $\mu(z)$ $=||K||^{2}+z\langle\phi_{\omega},H_{z}^{\prime}\phi_{\omega}\rangle+{\mathcal{O}}(|z|^{2})$
(3.24) $=||K||^{2}+z\langle\phi_{\omega},V\phi_{\omega}\rangle+{\mathcal{O}}(|z|^{2}).$
It therefore suffices to prove that
(3.25) $\langle\phi_{\omega},V\phi_{\omega}\rangle=\sum_{j\in{\mathbb{J}}}\omega^{j}V_{m(j)}.$
Towards that end, first recall that $V$, viewed as a multiplication operator, commutes with $g_{*}$ for all $g\in G$ by virtue of $V$ being $\Lambda$-invariant. Consequently, by expanding $\phi_{\omega}$ via (3.19), we compute that
(3.26) $\langle\phi_{\omega},V\phi_{\omega}\rangle$ $=\frac{1}{\sqrt{|G_{0}|}}\sum_{j\in{\mathbb{J}}}\omega^{-j}\left\langle e^{ig^{-j}K\cdot x},V\phi_{\omega}\right\rangle$
(3.27) $=\frac{1}{\sqrt{|G_{0}|}}\sum_{j\in{\mathbb{J}}}\omega^{-j}\left\langle e^{iK\cdot x},V(g_{*}^{j}\phi_{\omega})\right\rangle=\frac{1}{\sqrt{|G_{0}|}}\sum_{j\in{\mathbb{J}}}\left\langle e^{iK\cdot x},V\phi_{\omega}\right\rangle=\sqrt{|G_{0}|}\left\langle e^{iK\cdot x},V\phi_{\omega}\right\rangle$
(3.28) $=\sum_{j\in{\mathbb{J}}}\omega^{j}\left\langle e^{iK\cdot x},Ve^{ig^{-j}K\cdot x}\right\rangle=\sum_{j\in{\mathbb{J}}}\omega^{j}\left\langle e^{i(K-g^{-j}K)\cdot x},V\right\rangle=\sum_{j\in{\mathbb{J}}}\omega^{j}V_{m(j)}.$
Note that in (3.28), we have used the fact that $K-g^{-j}K=m(j)\cdot(k_{1},\ldots,k_{n})$. ∎
Proof of Lemma 3.3.
Since $z\in{\mathbb{R}}$ is assumed to be fixed, for simplicity of notation we suppress the dependence of ${\mathcal{E}}(z)$ on $z$, and denote this eigenspace simply by ${\mathcal{E}}$. We also note that the operators $H_{z}$ on $L^{2}_{K+\kappa}$ and $H_{z,\kappa}:=e^{i\kappa\cdot x}H_{z}e^{i\kappa\cdot x}$ on $L^{2}_{K}$ have the same spectrum. Indeed, if $\phi(x)\in L^{2}_{K}$, then $\psi(x):=e^{i\kappa\cdot x}\phi(x)\in L^{2}_{K+\kappa}$. Furthermore, $\psi$ is an $L^{2}_{K+\kappa}$ eigenvector of $H_{z}$ with eigenvalue $\mu$ if and only if
(3.29) $H_{z,\kappa}\phi(x)=e^{i\kappa\cdot x}H_{z}\psi(x)=\mu\phi(x),$
or, in other words, $\phi$ is an $L^{2}_{K}$ eigenvector of $H_{z,\kappa}$ with eigenvalue $\mu$. Therefore, to compute the $L^{2}_{K+\kappa}$-eigenvalues of $H_{z}$ near $\mu(z)$, we instead compute the $L^{2}_{K}$-eigenvalues of $H_{z,\kappa}$.
Write $H_{z,\kappa}$ as a $2\times 2$ block operator with respect to the decomposition $L^{2}_{K}={\mathcal{E}}\oplus{\mathcal{E}}^{\perp}$:
(3.30) $H_{z,\kappa}=\begin{pmatrix}[1.5]H_{z,\kappa}^{(11)}&H_{z,\kappa}^{(12)}\\ H_{z,\kappa}^{(21)}&H_{z,\kappa}^{(22)}\end{pmatrix}.$
Letting $\pi(z)^{\perp}=I-\pi(z)$, we compute that
(3.31) $H_{z,\kappa}^{(11)}$ $=\pi(z)H_{z,\kappa}\pi(z)=\pi(z)e^{i\kappa\cdot x}(-\Delta+zV(x))e^{i\kappa\cdot x}\pi(z)=\pi(z)(H_{z}-2i\kappa\cdot\nabla+|\kappa|^{2})\pi(z)$
(3.32) $=\mu(z)+M(z,\kappa)+{\mathcal{O}}(|\kappa|^{2}),$
(3.33) $H_{z,\kappa}^{(12)}$ $=\pi(z)(H_{z}-2i\kappa\cdot\nabla+|\kappa|^{2})\pi(z)^{\perp}=-\pi(z)(2i\kappa\cdot\nabla)\pi(z)^{\perp}+{\mathcal{O}}(|\kappa|^{2})={\mathcal{O}}(|\kappa|),$
(3.34) $H_{z,\kappa}^{(12)}$ $=\left(H_{z,\kappa}^{(12)}\right)^{*}={\mathcal{O}}(|\kappa|).$
Next, we claim that there exist $\varepsilon,\delta>0$ such that for $|\kappa|<\varepsilon$ and $\mu\in{\mathbb{B}}_{\delta}(\mu(z))$, $H^{(22)}_{z,\kappa}-\mu$ is invertible and its inverse is uniformly bounded in $\mu$. To prove this, observe that since $H_{z}$ has a compact resolvent and is self-adjoint due to our assumption that $z\in{\mathbb{R}}$, we can order the distinct eigenvalues of $H_{z}$ so that there exist eigenvalues $\mu_{-}<\mu(z)<\mu_{+}$ and the remaining eigenvalues of $H_{z}$ are all strictly farther away from $\mu(z)$. Consequently, if we let
(3.35) $\delta=\frac{1}{2}\operatorname{min}\big{\{}|\mu(z)-\mu_{-}|,|\mu(z)-\mu_{+}|\big{\}},$
it then follows that $H_{z}|_{{\mathcal{E}}^{\perp}}$ has no eigenvalues in ${\mathbb{B}}_{\delta}(\mu(z))$ since by construction $\mu(z)$ is not an eigenvalue of $H_{z}|_{{\mathcal{E}}^{\perp}}$. Therefore $H_{z}|_{{\mathcal{E}}^{\perp}}-\mu$ is invertible and its norm is bounded by $1/\delta$. In addition, the fact that $H_{z}$ is self-adjoint implies that $\operatorname{Im}(H_{z}|_{{\mathcal{E}}^{\perp}})\subset{\mathcal{E}}^{\perp}$, and as a result $(H_{z}-\mu)|_{{\mathcal{E}}^{\perp}}^{-1}$ is a well-defined operator from ${\mathcal{E}}^{\perp}$ to itself (and in fact is a bijection from ${\mathcal{E}}^{\perp}$ to $H^{2}_{K}\cap{\mathcal{E}}^{\perp}$ by elliptic regularity).
Thus, for all $\mu\in{\mathbb{B}}_{\delta}(\mu(z))$, we have
(3.36) $H^{(22)}_{z,\kappa}-\mu$ $=\pi(z)^{\perp}(H_{z}-\mu-2i\kappa\cdot\nabla+|\kappa|^{2})\pi(z)^{\perp}$
(3.37) $=(H_{z}-\mu)|_{{\mathcal{E}}^{\perp}}-\pi(z)^{\perp}(2i\kappa\cdot\nabla)\pi(z)^{\perp}+|\kappa|^{2}$
(3.38) $=(H_{z}-\mu)|_{{\mathcal{E}}^{\perp}}^{-1}\left(I-T(\mu,\kappa)\right),$
where
(3.39) $T(\mu,\kappa)=(H_{z}-\mu)|_{{\mathcal{E}}^{\perp}}^{-1}\left(\pi(z)^{\perp}(2i\kappa\cdot\nabla)\pi(z)^{\perp}+|\kappa|^{2}\right).$
Again using elliptic regularity, for each $\mu\in{\mathbb{B}}_{\delta}(\mu(z))$ there exists some $C_{\mu}>0$ such that $||T(\mu,\kappa)||\leq C_{\mu}|\kappa|$. In addition, since ${\mathbb{B}}_{\delta}(\mu(z))$ is contained in the resolvent set of $H_{z}$, $T(z,\kappa)$ is continuous in $\mu$ on this set, which together with the fact that ${\mathbb{B}}_{\delta}(\mu(z))$ is precompact, means that there exists some $C>0$, uniform in $\mu$, such that $||T(\mu,\kappa)||\leq C|\kappa|$. Therefore, if we set $\varepsilon=1/C$, it follows from a Neumann series argument such as the one following (2.38) that $H^{(22)}_{z,\kappa}-\mu$ is invertible and its inverse satisfies
(3.40) $(H^{(22)}_{z,\kappa}-\mu)^{-1}=(H_{z}-\mu)|_{{\mathcal{E}}^{\perp}}^{-1}+{\mathcal{O}}(|\kappa|),$
uniformly in $\mu$, for $\mu\in{\mathbb{B}}_{\delta}(\mu(z))$ and $|\kappa|<\varepsilon$.
Now pick some such $\kappa$, let $\lambda_{1}(z,\kappa),\ldots,\lambda_{r}(z,\kappa)$ denote the distinct eigenvalues of $M(z,\kappa)$, and fix some $j\in\{1,\ldots,r\}$. Viewing $M(z,\kappa)$ as a finite-dimensional operator on ${\mathcal{E}}$, note that $M(z,0)$ is the zero matrix, and consequently, after shrinking $\varepsilon$ if necessary, we may assume by continuity of eigenvalues $|\lambda_{j}(z,\kappa)|<\delta$. There then exists a simple, closed, positively-oriented contour ${\mathcal{C}}$ contained in ${\mathbb{B}}_{\delta}(\mu(z))$, such that ${\mathcal{C}}$ encloses $\mu(z)+\lambda_{j}(z,\kappa)$ and no other eigenvalue of $\mu(z)+M(z,\kappa)$. Furthermore, since ${\mathcal{C}}$ is compact and $((\mu(z)-\mu)+M(z,\kappa))^{-1}$ is continuous in $\mu$ for all $\mu\in{\mathcal{C}}$, there exists $M>0$ such that
(3.41) $||((\mu(z)-\mu)+M(z,\kappa))^{-1}||\leq M$
for all $\mu\in{\mathcal{C}}$.
We now want to use the Schur complement formula and Cauchy’s integral formula to relate the eigenvalues of $H_{z,\kappa}$ to those of $M(z,\kappa)$. Since $H^{(22)}_{z,\kappa}-\mu$ is invertible for all $\mu\in{\mathcal{B}}_{\delta}(\mu(z)$ and $|\kappa|<\varepsilon$, the Schur complement of the block $H^{(22)}_{z,\kappa}-\mu$ is well-defined for all such $\mu$ and $\kappa$ and is given by:
(3.42) $(H^{(11)}_{z,\kappa}-\mu)+H^{(12)}_{z,\kappa}(H^{(22)}_{z,\kappa}-\mu)^{-1}H^{(21)}_{z,\kappa}=(\mu(z)-\mu)+M(z,\kappa)+{\mathcal{O}}(|\kappa|^{2}),$
uniformly in $\mu$, where we have used the fact that $||(H^{(22)}_{z,\kappa}-\mu)^{-1}||\leq\delta+C|\kappa|$ (and this bound is also uniform in $\mu$). We then claim that the operator given by (3.42) is invertible and its inverse is uniformly bounded in $\mu$ for all $\mu\in{\mathcal{C}}$. To prove this claim, first observe that
(3.43) $(\mu(z)-\mu)+M(z,\kappa)+{\mathcal{O}}(|\kappa|^{2})$ $=((\mu(z)-\mu)+M(z,\kappa))\left(I+S(z,\kappa)\right),\quad\text{where }$
(3.44) $S(z,\kappa)$ $=((\mu(z)-\mu)+M(z,\kappa))^{-1}\cdot{\mathcal{O}}(|\kappa|^{2}).$
By (3.41), after increasing $M$ if necessary, $||S(z,\kappa)||\leq M|\kappa|^{2}$ for all $\mu\in{\mathcal{C}}$. Therefore, by replacing $\varepsilon$ with the minimum of itself and $1/\sqrt{M}$, it follows from another Neumann series argument that $(\mu(z)-\mu)+M(z,\kappa)+{\mathcal{O}}(|\kappa|^{2})$ is invertible and its inverse satisfies
(3.45) $\left(((\mu(z)-\mu)+M(z,\kappa)+{\mathcal{O}}(|\kappa|^{2})\right)^{-1}=((\mu(z)-\mu)+M(z,\kappa))^{-1}+{\mathcal{O}}(|\kappa|^{2})$
for all $\mu\in{\mathcal{C}}$ and $|\kappa|<\varepsilon$, where again the bound is uniform in $\mu$.
Thus, for all such $\mu$ and $\kappa$, we can write $(H_{z,\kappa}-\mu)^{-1}$ with respect to the decomposition $L^{2}_{K}={\mathcal{E}}\oplus{\mathcal{E}}^{\perp}$ as
(3.46) $(H_{z,\kappa}-\mu)^{-1}=\begin{pmatrix}[1.5]((\mu(z)-\mu)+M(z,\kappa))^{-1}+{\mathcal{O}}(|\kappa|^{2})&{\mathcal{O}}(|\kappa|)\\ {\mathcal{O}}(|\kappa|)&(H^{(22)}_{z,\kappa}-\mu)^{-1}+{\mathcal{O}}(|\kappa|^{2})\end{pmatrix},$
where all bounds are uniform in $\mu$. Consequently, by applying Cauchy’s integral formula and taking the trace, we get
[…]
(3.49) $=m_{j}\cdot(\mu(z)+\lambda_{j}(z,\kappa))+{\mathcal{O}}(|\kappa|^{2}),$
where $m_{j}$ is the multiplicity of $\lambda_{j}(z,\kappa)$ as an eigenvalue of $M(z,\kappa)$. In (3.49), we have used the fact that $\mu(H^{(22)}_{z,\kappa}-\mu)^{-1}$ is analytic in $\mu$ on ${\mathbb{B}}_{\delta}(\mu(z))$, and so its integral on ${\mathcal{C}}$ equals zero, and the fact that the integral of $\mu\cdot{\mathcal{O}}(|\kappa|^{2})$ is ${\mathcal{O}}(|\kappa|^{2})$, due to the ML inequality and since the bound is uniform in $\mu$. We therefore conclude that the eigenvalues of $H_{\kappa}$ near $\mu(K_{0})$ are approximately given by the eigenvalues of $\mu(K_{0})+M(\kappa)$, up to order $|\kappa|^{2}$. ∎
Proof of Lemma 3.4.
Let $m=\operatorname{dim}{\mathcal{E}}(z)$, which is constant by analyticity of $\pi(z)$ and [kato, Lemma I.4.10]. As a symmetric function of eigenvalues, the determinant of $M(z,\kappa)-\lambda$ can be expressed as a (universal) polynomial in the traces of its $m$ first powers. This means that $\operatorname{det}_{{\mathcal{E}}(z)}(M(z,\kappa)-\lambda)$ is polynomial (with coefficients independent of $z$, $\kappa$, and $\lambda$) in
(3.50) $\operatorname{Tr}_{{\mathcal{E}}(z)}\big{(}(M(z,\kappa)-\lambda)^{j}\big{)}=\operatorname{Tr}_{L^{2}_{K}}\big{(}(M(z,\kappa)-\lambda\pi(z))^{j}\big{)},\quad\text{ for }j=1,\ldots,m.$
The operator $M(z,\kappa)-\lambda\pi(z)$ is finite-rank and analytic in $z$, and hence its trace is analytic in $z$. Thus $\operatorname{det}_{{\mathcal{E}}(z)}\big{(}M(z,\kappa)-\lambda\big{)}$ is analytic in $z$. ∎
Proof of Lemma 3.5.
We start by looking at how the group action of $G_{0}$ interacts with the gradient of a function $f\in L^{2}_{K}$. Let $g\in G_{0}$; then
(3.51) $\nabla(g_{*}f)(x)=\nabla(f(g^{\top}x))=g(\nabla f)(g^{\top}x)=g(g_{*}\nabla f)(x).$
Multiplying the first and last of these expressions on the right by $g^{\top}$, we get
(3.52) $g_{*}\nabla f=g^{\top}\nabla(g_{*}f).$
Now let $\phi\in L^{2}_{K,\omega}$ and let $\psi\in L^{2}_{K,\widetilde{\omega}}$ for $\omega,\widetilde{\omega}\in{\mathbb{U}}$. To show that $\langle\phi,\nabla\psi\rangle$ is an eigenvector of $g^{j}$ with corresponding eigenvalue $\omega^{-j}\widetilde{\omega}^{j}$ for all $j\in{\mathbb{J}}$, we compute the following, using (3.52) and the fact that $(g^{j})^{\top}=g^{-j}$ since $G_{0}$ consists of orthogonal matrices:
(3.53) $g^{j}\langle\phi,\nabla\psi\rangle=g^{j}\langle g^{j}_{*}\phi,g^{j}_{*}\nabla\psi\rangle=\langle g^{j}_{*}\phi,\nabla(g^{j}_{*}\psi)\rangle=\overline{\omega^{j}}\widetilde{\omega}^{j}\langle\phi,\nabla\psi\rangle=\omega^{-j}\widetilde{\omega}^{j}\langle\phi,\nabla\psi\rangle.$
Lastly, to show that $\langle\phi,\nabla\phi\rangle=0$ when $K$ is a vertex of ${\mathcal{B}}$, note that (3.53) implies that $\langle\phi,\nabla\phi\rangle$ is an eigenvector of $g^{j}$ with eigenvalue 1 for all $j\in{\mathbb{J}}$. However, we claim that the only such vector is the zero vector. Suppose $k\in{\mathbb{R}}^{n}$ such that $gk$ for all $g\in G_{0}$. Then, since $K$ is assumed to be a vertex of ${\mathcal{B}}$, it necessarily must lie on at least $n$ hyperfaces of ${\mathcal{B}}$, and therefore, by Proposition A.1, there exist linearly independent lattice vectors $K_{1},\ldots,K_{n}\in\Lambda^{*}$ such that $K-K_{j}$ is also a vertex of ${\mathcal{B}}$ satisfying $||K-K_{j}||=||K||$, and therefore contained in the equivalence class $[K]$. By 3.3, for $j=1,\ldots,n$, there exists $h_{j}\in G_{0}$ (where the notation here is chosen so as to differentiate $h_{j}$ from the generator $g_{j}$) such that $h_{j}K=K-K_{j}$. Thus we get that, for all $j$,
(3.54) $k\cdot K_{j}=k\cdot(K-h_{j}K)=k\cdot K-h_{j}^{\top}k\cdot K=k\cdot K-k\cdot K=0.$
Since the set $\{K_{j}\}_{j=1}^{n}$ is linearly independent, it is a basis for ${\mathbb{R}}^{n}$, and it therefore follows that $k=0$. As a result, we conclude that $\langle\phi,\nabla\phi\rangle=0$. ∎
4. Schrödinger Operators Invariant Under Cubic Lattices
In this section, we focus on Schrödinger operators invariant under cubic lattices, which are lattices whose point groups are isomorphic to the octahedral group. Every such lattice is isometric, up to a dilation, to one of the three lattices generated by the bases listed in row 1 of Table 4; these lattices are called the simple cubic, body-centered cubic, and face-centered cubic, respectively. Using the general theory developed in §3, we prove that the dispersion surfaces of such Schrödinger operators generically have unusual dispersion surfaces near vertices of the Brillouin zone: Theorem 2.
4.1. Geometry of Cubic Lattices.
$\Lambda^{S}$ $\Lambda^{BC}$ $\Lambda^{FC}$
Basis $\begin{pmatrix}1\\ 0\\ 0\end{pmatrix},\begin{pmatrix}0\\ 1\\ 0\end{pmatrix},\begin{pmatrix}0\\ 0\\ 1\end{pmatrix}$ $\begin{pmatrix}1\\ 0\\ 0\end{pmatrix},\begin{pmatrix}0\\ 1\\ 0\end{pmatrix},\begin{pmatrix}1/2\\ 1/2\\ 1/2\end{pmatrix}$ $\begin{pmatrix}1/2\\ 1/2\\ 0\end{pmatrix},\begin{pmatrix}-1/2\\ 1/2\\ 0\end{pmatrix},\begin{pmatrix}0\\ -1/2\\ 1/2\end{pmatrix}$
Dual Basis $\begin{pmatrix}2\pi\\ 0\\ 0\end{pmatrix},\begin{pmatrix}0\\ 2\pi\\ 0\end{pmatrix},\begin{pmatrix}0\\ 0\\ 2\pi\end{pmatrix}$ $\begin{pmatrix}2\pi\\ 0\\ -2\pi\end{pmatrix},\begin{pmatrix}0\\ 2\pi\\ -2\pi\end{pmatrix},\begin{pmatrix}0\\ 0\\ 4\pi\end{pmatrix}$ $\begin{pmatrix}2\pi\\ 2\pi\\ 2\pi\end{pmatrix},\begin{pmatrix}-2\pi\\ 2\pi\\ 2\pi\end{pmatrix},\begin{pmatrix}0\\ 0\\ 4\pi\end{pmatrix}$
${\mathcal{B}}$ Refer to caption Refer to caption Refer to caption
$K$ $\begin{pmatrix}\pi\\ \pi\\ \pi\end{pmatrix}$ $\begin{pmatrix}\pi\\ \pi\\ \pi\end{pmatrix}$ $\begin{pmatrix}0\\ 0\\ 2\pi\end{pmatrix}$ $\begin{pmatrix}0\\ 2\pi\\ \pi\end{pmatrix}$
$m$ 8 4 6 4
$G_{0}$ $\langle f_{1},f_{2},f_{3}\rangle$ $\langle f_{13},f_{23}\rangle$ $\langle r,f\rangle$ $\langle s_{0}\rangle$
${\mathbb{U}}$ $U_{2}^{3}$ $U_{2}^{2}$ $U_{3}\times U_{2}$ $U_{4}$
Figure 4. Geometry of the cubic lattices.
Let $\Lambda^{S}$, $\Lambda^{BC}$, and $\Lambda^{FC}$ denote the simple cubic, body-centered cubic, and face-centered cubic lattice, respectively, which are generated by the bases given in row 1 of Table 4. We then give spectral results for $-\Delta$ seen as a $\Lambda^{S}$, $\Lambda^{BC}$, and $\Lambda^{FC}$ invariant operator on $L^{2}_{K}$, where $K$ is a vertex of the corresponding Brillouin zone ${\mathcal{B}}$, as these points exhibit a high degree of symmetry (see Proposition A.1). In addition, as noted in §3.2, it suffices to consider vertices $K$ which have distinct orbits under the action of the point group $G$.
As previously discussed, the point group of the lattices $\Lambda^{S},\Lambda^{BC}$ and $\Lambda^{FC}$ is the octahedral group, which we denote by $G$, and which is generated by the three matrices
(4.1) $r:=\begin{pmatrix}0&0&1\\ 1&0&0\\ 0&1&0\end{pmatrix},\quad s:=\begin{pmatrix}0&1&0\\ 1&0&0\\ 0&0&-1\end{pmatrix},\quad\text{and}\quad f:=\begin{pmatrix}-1&0&0\\ 0&-1&0\\ 0&0&-1\end{pmatrix}.$
We shall also later need the following elements of $G$:
(4.2) $f_{1}:=\begin{pmatrix}-1&0&0\\ 0&1&0\\ 0&0&1\end{pmatrix},\quad f_{2}:=\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&1\end{pmatrix},\quad f_{3}:=\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&-1\end{pmatrix},$
(4.3) $f_{12}:=\begin{pmatrix}-1&0&0\\ 0&-1&0\\ 0&0&1\end{pmatrix},\quad f_{13}:=\begin{pmatrix}-1&0&0\\ 0&1&0\\ 0&0&-1\end{pmatrix},\quad f_{23}:=\begin{pmatrix}1&0&0\\ 0&-1&0\\ 0&0&-1\end{pmatrix},$
[…]
In Table 4, in addition to a basis $v_{1},v_{2},v_{3}$, we list for each of the lattices $\Lambda^{S}$, $\Lambda^{BC}$, and $\Lambda^{FC}$:
  • The corresponding dual basis $k_{1},k_{2},k_{3}$ (which by definition satisfies $v_{j}\cdot k_{\ell}=2\pi\delta_{j\ell}$ for $j,\ell=1,2,3$);
  • A picture of the Brillouin zone ${\mathcal{B}}$, where the vertices are colored and sized in reference to the vertices listed in the following row. Specifically, given a vertex $K$, the set $[K]$ (defined in (3.13)) consists of vertices of ${\mathcal{B}}$ by Proposition A.2, which are colored the same and have larger dots. Those vertices which lie in the same orbit under $G$ but are not in $[K]$ are colored the same but have smaller dots;
  • Vertices $K$ of the Brillouin zone, corresponding to distinct orbits under the action of $G$.
  • The multiplicity $m$ of the $L^{2}_{K}$-eigenvalue $||K||^{2}$ of $-\Delta$, equal to the cardinality of the set $[K]$;
  • An abelian subgroup $G_{0}$ of $G$, expressed in terms of its generators, which together with the vertex $K$ satisfy 3.3;
  • The corresponding group ${\mathbb{U}}$ consisting of tuples of roots of unity, as defined in (3.15).
4.2. Proof Outline for Theorem 2
In each of the following four sections, we shall prove Theorem 2 for one of the three cubic lattices together with one of the vertices $K$ listed in row 4 of Table 4 by using the lemmas stated in §3.4. Each of these proofs will require the same three steps, which we now outline.
(1) Upper bound on multiplicity: Let $\Lambda$ be one of the three cubic lattices listed in Table 4, let $K$ be one of the listed vertices for this lattice, and let $(k_{1},k_{2},k_{3})$, ${\mathcal{B}}$, $m$, $G_{0}$, and ${\mathbb{U}}$ be the objects listed in the column corresponding to this vertex. We also let $V$ be a $\Lambda$-invariant potential and let $H_{z}=-\Delta+zV$. Then for each $\omega\in{\mathbb{U}}$, Lemma 3.2 describes how the multiplicity $m$, $L^{2}_{K}$-eigenvalue $||K||^{2}$ of $H_{0}=-\Delta$ splits as $z$ increases into simple $L^{2}_{K,\omega}$-eigenvalues given by:
(4.5) $\mu_{\omega}(z)=||K||^{2}+z\cdot\sum_{j\in{\mathbb{J}}}\omega^{j}V_{m(j)}+{\mathcal{O}}(|z|^{2}),$
In particular, if $\omega,\widetilde{\omega}\in{\mathbb{U}}$ are such that $\mu^{\prime}_{\omega}(0),\mu^{\prime}_{\widetilde{\omega}}(0)$ are distinct, then the eigenvalues $\mu_{\omega}(z),\mu_{\widetilde{\omega}}(z)$ clearly split. This test provides an upper bound on possible multiplicities of $\mu_{\omega}(z)$, viewed as an $L^{2}_{K}$-eigenvalue.
(2) Lower bound on multiplicity: Our argument in step (1) is inconclusive when $\mu^{\prime}_{\omega}(0)=\mu^{\prime}_{\widetilde{\omega}}(0)$, and so in this case we provide a lower bound on the splitting multiplicities using a symmetry argument. Note that the multiplicity of $\mu_{\omega}(z)$ as an $L^{2}_{K}$-eigenvalue is at least one, so it suffices to prove a lower bound on the multiplicity of $\mu_{\omega}(z)$ for those $\omega\in{\mathbb{U}}$ such that the upper bound computed in (1) is strictly greater than 1. This argument will typically rely on the existence of symmetries $S$ of $H_{z}$ such that
(4.6) $S\big{(}L^{2}_{K,\omega}\big{)}=L^{2}_{K,\widetilde{\omega}}.$
This implies that $H_{z}$ on $L^{2}_{K,\lambda}$ and $L^{2}_{K,\lambda}$ are conjugated, hence isospectral: $\mu_{\omega}(z)=\mu_{\widetilde{\omega}}(z)$. In each case, this will provide a lower bound on the multiplicity of $\mu_{\omega}(z)$ as an $L^{2}_{K}$-eigenvalue equaling the upper bound computed in step (1), and thus we deduce that $\mu_{\omega}(z)$ has constant multiplicity, which we denote by $m_{\omega}$, for sufficiently small $z$. It then follows that $\mu_{\omega}(z)$ satisfies 1.1, and so we can apply Theorem 1 to conclude that $H_{z}$ has an $L^{2}_{K}$-eigenvalue $\mu_{\omega}(z)$ which for generic $z\in{\mathbb{R}}$ has multiplicity $m_{\omega}$ and whose eigenprojector $\pi_{\omega}(z)$ is analytic on a open, connected neighborhood $U$ of the real line away from a discrete set.
(3) Computation of the characteristic polynomial: Let
(4.7) ${\mathbb{U}}_{\omega}=\{\widetilde{\omega}\in{\mathbb{U}}\;:\;\mu^{\prime}_{\widetilde{\omega}}(0)=\mu^{\prime}_{\omega}(0)\},$
so that $m_{\omega}=|{\mathbb{U}}_{\omega}|$ and $\mu_{\omega}=\mu_{\widetilde{\omega}}$ for all $\widetilde{\omega}\in{\mathbb{U}}_{\omega}$ by steps (1) and (2). Then for all such $\widetilde{\omega}$ and sufficiently small $z$,
(4.8) $\big{(}H_{z}-\mu_{\omega}(z)\pi_{\omega}(z)\big{)}|_{L^{2}_{K,\widetilde{\omega}}}=0,$
and by analyticity this must hold for all $z\in U$. Therefore $\mu_{\omega}(z)$ is a simple $L^{2}_{K,\widetilde{\omega}}$-eigenvalue for all such $z$ and $\widetilde{\omega}\in{\mathbb{U}}_{\omega}$, and as a result there exists a basis $\phi_{1},\ldots,\phi_{m_{\omega}}$, normalized to have $L^{2}_{K}$-norm 1, of the eigenspace corresponding to $\mu_{\omega}(z)$ consisting of precisely one vector from $L^{2}_{K,\widetilde{\omega}}$ for each $\widetilde{\omega}\in{\mathbb{U}}_{\omega}$.
Now let $z\in U\cap{\mathbb{R}}$; Lemma 3.3 then describes the structure of the dispersion surfaces corresponding to $\mu_{\omega}(z)$ near the vertex $K$:
(4.9) $\mu_{\omega}(z,\kappa)=\mu(z)+\lambda_{j}(z,\kappa)+{\mathcal{O}}(|\kappa|^{2}),$
where $\lambda_{j}(z,\kappa)$ are the eigenvalues of $M(z,\kappa)=-\pi_{\omega}(z)(2i\kappa\cdot\nabla)\pi_{\omega}(z)$ on the eigenspace corresponding to $\mu(z)$. Using the basis $\phi_{1},\ldots,\phi_{m_{\omega}}$, we can then compute the entries of $M(z,\kappa)$ with respect to this basis using Lemma 3.5. In particular, this lemma tells us that the diagonal entries of $M(z,\kappa)$ are all zero, and we only need to compute the entries above the diagonal since $M(z,\kappa)$ is Hermitian.
Lastly, once we have an explicit expression for $M(z,\kappa)$, we can compute its characteristic polynomial. We then finish by checking which of the coefficients of this polynomial are nonzero for $z$ sufficiently small, which by Lemma 3.4 will then imply that these coefficients remain nonzero for all $z\in U\cap{\mathbb{R}}$ away from a discrete set in $U$ by analyticity. To perform this computation, we will typically use the fact that a normalized eigenvector $\phi_{\omega}(x;z)$ corresponding to the $L^{2}_{K,\omega}$-eigenvalue $\mu_{\omega}(z)$ satisfies
(4.10) $\phi_{\omega}(x;z)=\phi_{\omega}(x)+{\mathcal{O}}(|z|),$
where $\phi_{\omega}(x)$ is the normalized eigenvector corresponding to $\mu_{\omega}(0)$ given by (3.19). This follows from the observation that $\pi_{\omega}(z)\phi_{\omega}$ is an eigenvector corresponding to $\mu_{\omega}(z)$ for $z$ sufficiently small, and the fact that $\pi_{\omega}(z)=\pi_{\omega}(0)+{\mathcal{O}}(|z|)$ by a Neumann series argument. We then conclude that the coefficients of the characteristic polynomial are nonzero for all $z\in{\mathbb{R}}$ except for a closed and countable set.
4.3. Proof of Theorem 2 for the Simple Cubic
Let $\Lambda=\Lambda^{S}$, and let $(k_{1},k_{2},k_{3})$, ${\mathcal{B}}$, $K$, $m$, $G_{0}$, and ${\mathbb{U}}$ be the objects listed in corresponding column (i.e. the first column) of Table 4. We also let $V$ be a $\Lambda$-invariant potential and let $H_{z}=-\Delta+zV$. Lastly we will need the group elements $r,s,f$ defined in (4.1) and $f_{1},f_{2},f_{3}$ defined in (4.2).
(1) Upper bound on multiplicity: For each $\omega\in{\mathbb{U}}$, $\mu^{\prime}_{\omega}(0)$ is given by:
(4.11) $\mu^{\prime}_{\omega}=\sum_{j\in{\mathbb{J}}}\omega^{j}V_{m(j)}.$
A quick computation shows that, for $j=1,2,3$,
(4.12) $f_{j}^{-1}K=K-k_{j}.$
It follows from the definition of $m(j)$ (given by (3.17)) that $m(j)=-j$. Moreover, since $V$ is even (as noted in §3.2), it follows that $V_{-j}=V_{j}$. Thus we have the formula
(4.13) $\mu^{\prime}_{\omega}=\sum_{j\in{\mathbb{J}}}\omega^{j}V_{j}.$
In addition, observe that $V$ is invariant under $r$, which permutes the coordinate axes. Consequently, we also have the identities
(4.14) $V_{1,0,0}=V_{0,1,0}=V_{0,0,1}\quad\text{ and }\quad V_{1,1,0}=V_{1,0,1}=V_{0,1,1}.$
It follows that we can rewrite $\mu^{\prime}_{\omega}(0)$ as
(4.15) $\mu^{\prime}_{\omega}(0)=V_{0,0,0}+(\omega_{1}+\omega_{2}+\omega_{3})V_{1,0,0}+(\omega_{2}\omega_{3}+\omega_{1}\omega_{3}+\omega_{1}\omega_{2})V_{1,1,0}+\omega_{1}\omega_{2}\omega_{3}V_{1,1,1}.$
We then plug $\omega$ into this formula for each $\omega\in{\mathbb{U}}$, which gives us the following:
(4.16) $\mu^{\prime}_{1,1,1}(0)$ $=V_{0,0,0}+3V_{1,0,0}+3V_{1,1,0}+V_{1,1,1},$
(4.17) $\mu^{\prime}_{-1,1,1}(0)=\mu^{\prime}_{1,-1,1}(0)=\mu^{\prime}_{1,1,-1}(0)$ $=V_{0,0,0}+V_{1,0,0}-V_{1,1,0}-V_{1,1,1},$
(4.18) $\mu^{\prime}_{1,-1,-1}(0)=\mu^{\prime}_{-1,1,-1}(0)=\mu^{\prime}_{-1,-1,1}(0)$ $=V_{0,0,0}-V_{1,0,0}-V_{1,1,0}+V_{1,1,1},$
(4.19) $\mu^{\prime}_{-1,-1,-1}(0)$ $=V_{0,0,0}-3V_{1,0,0}+3V_{1,1,0}-V_{1,1,1}.$
Note that the set where the right-hand sides of any pair of the above 4 equations are equal describes a hyperplane. Consequently, the set where the right-hand sides of the above four equations fail to be distinct is a union of six hyperplanes. It follows that for $V$ in a set of codimension 1, the eigenvalue $||K||^{2}$ of $-\Delta$ splits into at least two simple eigenvalues and two eigenvalues of multiplicity at most three.
(2) Lower bound on multiplicity: Observe that $f_{1}r=rf_{3}$, $f_{2}r=rf_{1}$, and $f_{3}r=rf_{2}$. As a result, if $\phi\in L^{2}_{K,(-1,1,1)}$ is an eigenvector of $H_{z}$, then $r_{*}\phi$ is also an eigenvector of $H_{z}$ with the same eigenvalue and
(4.20) $(f_{1})_{*}(r_{*}\phi)$ $=r_{*}(f_{3})_{*}\phi=r_{*}\phi$
(4.21) $(f_{2})_{*}(r_{*}\phi)$ $=r_{*}(f_{1})_{*}\phi=-r_{*}\phi$
(4.22) $(f_{3})_{*}(r_{*}\phi)$ $=r_{*}(f_{2})_{*}\phi=r_{*}\phi.$
Hence, $r_{*}\phi\in L^{2}_{K_{S},(1,-1,1)}$, and an identical computation shows that $r^{2}_{*}\phi$ is an eigenvector of $H_{z}$ as well, but in $L^{2}_{K,(1,1,-1)}$. Therefore $\mu_{(-1,1,1)}(z)$ is an $L^{2}_{K}$-eigenvalue with multiplicity at least 3, which together with step (1) implies that its multiplicity is exactly 3. The same argument applied to an eigenvector $\phi$ of $H_{z}$ in $L^{2}_{K,(1,-1,-1)}$ shows that $\mu_{(-1,1,1)}(z)$ is an $L^{2}_{K}$-eigenvalue with multiplicity 3 as well. Therefore $H_{z}$ has two triple $L^{2}_{K}$-eigenvalues for all $z\in{\mathbb{R}}$ away from a discrete set, and the corresponding eigenprojectors are analytic on some open set $U$.
(3) Computation of the characteristic polynomial: Fix some $z\in U\cap{\mathbb{R}}$, and let $\phi_{1}\in L^{2}_{K,(-1,1,1)},\phi_{2}\in L^{2}_{K,(1,-1,1)}$ and $\phi_{3}\in L^{2}_{K,(1,1,-1)}$ be normalized eigenvectors for the eigenvalue $\mu_{(-1,1,1)}(z)$ of $H_{z}$. The entries of $M(z,\kappa)$ with respect to this basis are given by $-2i\kappa\cdot\langle\phi_{j},\nabla\phi_{\ell}\rangle$. For $j\not=\ell$, $\langle\phi_{j},\nabla\phi_{\ell}\rangle$ is an eigenvector of both $f_{j}$ and $f_{\ell}$ with eigenvalue -1 by Lemma 3.5, and thus it lies in ${\mathbb{C}}e_{j}\cap{\mathbb{C}}e_{\ell}=\{0\}$. It follows that $M(z,\kappa)=0$ for all $z$ and $\kappa$, and thus we conclude that
(4.23) $\mu_{(-1,1,1)}(z,\kappa)=\mu_{(-1,1,1)}(z)+{\mathcal{O}}(|\kappa|^{2}).$
By Definition 1 this means that $(K,\mu_{(-1,1,1)}(z))$ is a 3-fold quadratic point. The exact same argument shows that $(K,\mu_{(-1,1,1)}(z))$ is a 3-fold quadratic point as well. This completes the proof of Theorem 2 when $\Lambda$ is a simple cubic lattice.
4.4. Proof of Theorem 2 for the Body-Centered Cubic at $K=(\pi,\pi,\pi)$
Let $\Lambda=\Lambda^{BC}$, let $K=(\pi,\pi,\pi)$ and let $(k_{1},k_{2},k_{3})$, ${\mathcal{B}}$, $m$, $G_{0}$, and ${\mathbb{U}}$ be the objects listed in the corresponding column of Table 4 (i.e. the second of the three columns for the first three rows and the second of the four columns for the remaining rows). We also let $V$ be a $\Lambda$-invariant potential and let $H_{z}=-\Delta+zV$. Lastly we will again need the generators $r,s,f$, and also the group elements $f_{12},f_{13},f_{23}$ defined in (4.3).
(1) Upper bound on multiplicity: Just as we did in §4.3, we start by computing relations among the Fourier coefficients $V_{m(j)}$ for $j\in{\mathbb{J}}$. To begin, we compute that
(4.24) $f_{13}^{-1}K$ $=K-k_{1}-k_{3},$
(4.25) $f_{23}^{-1}K$ $=K-k_{2}-k_{3},$
(4.26) $f_{12}^{-1}K$ $=K-k_{1}-k_{2}-k_{3}.$
Again using the fact that $V$ is even, it follows that for $\omega\in{\mathbb{U}}$,
(4.27) $\mu^{\prime}_{\omega}(0)=V_{0,0,0}+\omega_{1}V_{1,0,1}+\omega_{2}V_{0,1,1}+\omega_{1}\omega_{2}V_{1,1,1}.$
Also note that $V$ being invariant under $r$ implies $V_{1,0,1}=V_{0,1,1}$. In addition, $V_{1,1,1}=V_{1,0,0}$ since
(4.28) $V_{k_{1}+k_{3}}=r_{*}V_{k_{1}+k_{3}}=V_{r^{\top}(k_{1}+k_{3})}=V_{k_{2}+k_{3}}.$
Therefore
(4.29) $\mu^{\prime}_{\omega}(0)=V_{0,0,0}+(\omega_{1}+\omega_{2}+\omega_{1}\omega_{2})V_{1,1,1}.$
We then plug $\omega$ into this formula for each $\omega\in{\mathbb{U}}$ to obtain:
(4.30) $\mu^{\prime}_{1,1}(0)$ $=V_{0,0,0}+3V_{1,1,1},$
(4.31) $\mu^{\prime}_{-1,1}(0)=\mu^{\prime}_{1,-1}(0)=\mu^{\prime}_{-1,-1}(0)$ $=V_{0,0,0}-V_{1,1,1}.$
The set where the right-hand sides of the above two equations fail to be distinct is a single hyperplane. Therefore we again conclude that for $V$ in a set of codimension 1, the eigenvalue $||K||^{2}$ of $-\Delta$ splits into at least a simple eigenvalue and an eigenvalue of multiplicity at most three.
(2) Lower bound on multiplicity: Just as did in step (2) of Section 4.2, observe that $f_{13}r=rf_{23}$ and $f_{23}r=rf_{12}$. As a result, if $\phi$ is an eigenvector of $H_{z}$ in $L^{2}_{K,(-1,1)}$, then $r_{*}\phi$ and $r^{2}_{*}\phi$ are again eigenvectors of $H_{z}$ with the same eigenvalue on $L^{2}_{K_{BC},(1,-1)}$ and $L^{2}_{K_{BC},(-1,-1)}$, respectively. Therefore $H_{z}$ has a triple $L^{2}_{K}$-eigenvalue for all $z\in{\mathbb{R}}$ away from a discrete set, and the corresponding eigenprojector is analytic on some open set $U$.
(3) Computation of the characteristic polynomial: Fix some $z\in U\cap{\mathbb{R}}$, and let $\phi_{1}\in L^{2}_{K,(-1,1)}$ be a normalized eigenvector of $H_{z}$ for the eigenvalue $\mu_{(-1,1)}(z)$. Then, as we saw in step (2), $\phi_{2}:=r_{*}\phi\in L^{2}_{K,(1,-1)}$ and $\phi_{3}:=r_{*}^{2}\phi\in L^{2}_{K,(-1,-1)}$ are eigenvectors of $H_{z}$ with eigenvalue $\mu_{(-1,1)}(z)$ as well, and thus form a basis for the corresponding eigenspace. The entries of $M(z,\kappa)$ with respect to this basis are given by $-2i\kappa\cdot\langle\phi_{j},\nabla\phi_{\ell}\rangle$, and we also note that $\langle\phi_{2},\nabla\phi_{3}\rangle=\langle r_{*}\phi_{1},\nabla(r_{*}\phi_{1})\rangle=r\langle\phi_{1},\nabla\phi_{2}\rangle$. Therefore, by Lemma 3.5, the entries of $M(z,\kappa)$ are entirely determined by $\langle\phi_{1},\nabla\phi_{2}\rangle$ and $\langle\phi_{1},\nabla\phi_{3}\rangle$.
To compute these, note that $f_{12}=f_{13}f_{23}$, and so again by Lemma 3.5,
(4.32) $f_{12}\langle\phi_{1},\nabla\phi_{2}\rangle=(-1)^{2}\langle\phi_{1},\nabla\phi_{2}\rangle=\langle\phi_{1},\nabla\phi_{2}\rangle.$
Hence, $\langle\phi_{1},\nabla\phi_{2}\rangle$ is an eigenvector of $f_{12}$ with eigenvalue 1, and is therefore of the form $\alpha e_{3}$ for some $\alpha\in{\mathbb{C}}$. An identical argument applied to $\langle\phi_{1},\nabla\phi_{3}\rangle$ and the element $f_{13}\in G$ implies that $\langle\phi_{1},\nabla\phi_{3}\rangle=\beta e_{2}$ for some $\beta\in{\mathbb{C}}$.
Note that $\alpha=-\overline{\beta}$:
(4.33) $\alpha=e_{3}\cdot\langle\phi_{1},\nabla\phi_{2}\rangle$ $=re_{2}\cdot\langle\phi_{1},\nabla\phi_{2}\rangle=e_{2}\cdot r^{2}_{*}\langle r^{2}_{*}\phi_{1},r^{2}_{*}\nabla\phi_{2}\rangle$
(4.34) $=e_{2}\cdot\langle\phi_{3},\nabla\phi_{1}\rangle=-e_{2}\cdot\overline{\langle\phi_{1},\nabla\phi_{3}\rangle}=-\overline{\beta}.$
Thus, with respect to the basis $\phi_{1},\phi_{2},\phi_{3}$, $M(z,\kappa)$ is given by:
(4.35) $M(z,\kappa)=-2i\begin{pmatrix}[1.5]0&\alpha\kappa_{3}&-\overline{\alpha}\kappa_{2}\\ -\overline{\alpha}\kappa_{3}&0&\alpha\kappa_{1}\\ \alpha\kappa_{2}&-\overline{\alpha}\kappa_{1}&0\\ \end{pmatrix}.$
A quick computation then gives the characteristic polynomial of $M(z,\kappa)$ (as a polynomial in $\mu$):
(4.36) $\mu^{3}-4|\alpha|^{2}||\kappa||^{2}\mu+16\operatorname{Im}(\alpha^{3})\kappa_{1}\kappa_{2}\kappa_{3}.$
By Lemma 3.4, the coefficients $|\alpha|^{2}$ and $\operatorname{Im}(\alpha^{3})$ are analytic in $z$, and therefore will be nonzero away from a discrete set if they are nonzero for $z$ sufficiently small. However, by Lemma 3.1 we can assume that, for $z$ sufficiently small, $\phi_{1},\phi_{2}$ are given by by:
(4.37) $\phi_{1}(x;z)$ $=e^{iK\cdot x}-e^{if_{13}K\cdot x}+e^{if_{23}K\cdot x}-e^{if_{12}K\cdot x}+{\mathcal{O}}(|z|),$
(4.38) $\phi_{2}(x;z)$ $=e^{iK\cdot x}+e^{if_{13}K\cdot x}-e^{if_{23}K\cdot x}-e^{if_{12}K\cdot x}+{\mathcal{O}}(|z|).$
It follows that, for $z$ small,
(4.39) $\alpha=e_{3}\cdot\langle\phi_{1},\nabla\phi_{2}\rangle$ $=\frac{i}{4}e_{3}\cdot\left(K-f_{13}K-f_{13}K+f_{12}K\right)+{\mathcal{O}}(|z|^{2})=\pi i+{\mathcal{O}}(|z|^{2}).$
Therefore both $|\alpha|^{2}$ and $\operatorname{Im}(\alpha^{3})$ are nonzero for $z$ sufficiently small, and thus remain nonzero for all $z\in U$ away from a discrete set. It follows that they are nonzero on $U\cap{\mathbb{R}}$ away from a discrete set, and thus are nonzero on ${\mathbb{R}}$ away from a closed and countable set. By Definition 1, we conclude that $(K,\mu_{(-1,1)}(z))$ is a 3-fold Weyl point.
4.5. Proof of Theorem 2 for the Body-Centered Cubic at $K=(0,0,\pi)$
Let $\Lambda=\Lambda^{BC}$, let $K=(0,0,\pi)$ and let $(k_{1},k_{2},k_{3})$, ${\mathcal{B}}$, $m$, $G_{0}$, and ${\mathbb{U}}$ be the objects listed in the corresponding column of Table 4 (i.e. the second of the three columns for the first three rows and the third of the four columns for the remaining rows). We also let $V$ be a $\Lambda$-invariant potential and let $H_{z}=-\Delta+zV$. Lastly we will need the group element $s_{0}$ defined in (4.4).
(1) Upper bound on multiplicity: We start by computing relations among the Fourier coefficients $V_{m(j)}$ for $j\in{\mathbb{J}}$. In particular, we compute that
(4.40) $r^{-1}K$ $=K+k_{2},$ $r^{-1}fK$ $=K-k_{2}-k_{3},$
(4.41) $r^{-2}K$ $=K+k_{1},$ $r^{-2}fK$ $=K-k_{1}-k_{3}$
(4.42) $f^{-1}K$ $=K-k_{3}.$
Again using the fact that $V$ is even, it follows that for $\omega\in{\mathbb{U}}$,
(4.43) $\mu^{\prime}_{\omega}(0)=V_{0,0,0}+\omega_{1}V_{0,1,0}+\omega_{1}^{2}V_{1,0,0}+\omega_{2}V_{0,0,1}+\omega_{1}\omega_{2}V_{0,1,1}+\omega_{1}^{2}\omega_{2}V_{1,0,1}.$
We also have that
(4.44) $V_{k_{1}}=(f_{1})_{*}V_{k_{1}}=V_{f_{1}^{\top}k_{1}}=V_{k_{1}+k_{3}},$
which tells us that $V_{1,0,0}=V_{1,0,1}$. Furthermore, since $V$ is invariant under $r$, we obtain $V_{1,0,0}=V_{0,1,0}=V_{1,0,1}=V_{0,1,1}$. Thus we can rewrite $\mu^{\prime}_{\omega}(0)$ as
(4.45) $\mu^{\prime}_{\omega}(0)=V_{0,0,0}+\big{(}\omega_{1}+\omega_{1}^{2}+\omega_{1}\omega_{2}+\omega_{1}^{2}\omega_{2}\big{)}V_{1,0,0}+\omega_{2}V_{0,0,1}.$
We then plug $\omega$ into this formula for each $\omega\in{\mathbb{U}}$ to obtain:
(4.46) $\mu^{\prime}_{1,1}(0)$ $=V_{0,0,0}+4V_{1,0,0}+V_{0,0,1},$
(4.47) $\mu^{\prime}_{\zeta_{3},1}(0)=\mu^{\prime}_{\overline{\zeta_{3}},1}(0)$ $=V_{0,0,0}-2V_{1,0,0}+V_{0,0,1},$
(4.48) $\mu^{\prime}_{1,-1}(0)=\mu^{\prime}_{\zeta_{3},-1}(0)=\mu^{\prime}_{\overline{\zeta_{3}},-1}(0)$ $=V_{0,0,0}-V_{0,0,1}.$
The set where the right-hand sides of the above three equations fail to be distinct is a union of three hyperplanes. Therefore we conclude that for $V$ in a set of codimension 1, the eigenvalue $||K||^{2}$ of $-\Delta$ splits into at least a simple eigenvalue, an eigenvalue of multiplicity at most two, and an eigenvalue of multiplicity at most three.
(2) Lower bound on multiplicity: Let $T$ be the conjugate-parity operator: $Tf(x)=\overline{f(-x)}$, and let $\phi_{1}$ be normalized eigenvector of $H_{z}$ in $L^{2}_{K,(\zeta_{3},1)}$ for $\mu_{(\zeta_{3},1)}(z)$. For $z\in{\mathbb{R}}$, $\phi_{2}=T\phi_{1}\in L^{2}_{K}$ is also an eigenvector of $H_{z}$ for the same eigenvalue since $V$ is even and real. In addition, observe that
(4.49) $r_{*}\phi_{2}(x)$ $=\overline{\phi_{1}(-r^{\top}x)}=\overline{\zeta_{3}\phi_{1}(-x)}=\overline{\zeta_{3}}\phi_{2}(x)$
(4.50) $f_{*}\phi_{2}(x)$ $=\overline{\phi_{1}(-f^{\top}x)}=\overline{\phi_{1}(-x)}=\phi_{2}(x).$
Therefore $\phi_{2}\in L^{2}_{K,(\overline{\zeta_{3}},1)}$.
To give a lower bound on the multiplicity of $\mu_{(1,-1)}(z)$, let $L^{2}_{K,-1}=\operatorname{ker}_{L^{2}_{K}}(f_{*}+1)$, i.e. the space of odd functions in $L^{2}_{K}$. By construction of the subspaces $L^{2}_{K,\omega}$, it follows that
(4.51) $L^{2}_{K,-1}=L^{2}_{K,(1,-1)}\oplus L^{2}_{K,(\zeta_{3},-1)}\oplus L^{2}_{K,(\overline{\zeta_{3}},-1)}.$
Also note that $s_{0}K=K-k_{3}$, and so $s_{0}$ is $K$-invariant, which together with the fact that $s_{0}$ commutes with $f$, implies that $(s_{0})_{*}$ is a well-defined operator on $L^{2}_{K,-1}$.
Now let
(4.52) $\psi_{1}(x)$ $=\sin(2\pi x_{1})+i\sin(2\pi x_{2}),$
(4.53) $\psi_{2}(x)$ $=\sin(2\pi x_{1})-i\sin(2\pi x_{2}),$
(4.54) $\psi_{3}(x)$ $=\sqrt{2}\sin(2\pi x_{3}).$
A quick computation confirms that $\psi_{j}\in L^{2}_{K,-1}$, $||\psi_{j}||=1$, and $-\Delta\psi_{j}=(2\pi)^{2}\psi_{j}=||K||^{2}\psi_{j}$ for $j=1,2,3$. In addition, note that $\sigma((s_{0})_{*})=U_{4}$, the fourth roots of unity. If we let ${\mathcal{E}}_{\omega}=L^{2}_{K,-1}\cap\operatorname{ker}_{L^{2}_{K,-1}}((s_{0})_{*}-\omega)$ for $\omega\in U_{4}$, then $\psi_{1}\in{\mathcal{E}}_{-i}$, $\psi_{2}\in{\mathcal{E}}_{i}$ and $\phi_{3}\in{\mathcal{E}}_{-1}$. Therefore $||K||^{2}$ is a simple eigenvalue of $-\Delta$ on ${\mathcal{E}}_{\omega}$ for $\omega\in\{-i,i,-1\}$, and so by Corollary 2.1 it follows that, for sufficiently small $z\in{\mathbb{R}}$, there is a unique eigenvalue $\lambda(z)$ of $H_{z}$ on ${\mathcal{E}}_{-i}$ satisfying $\lambda(z)=||K||^{2}+{\mathcal{O}}(|z|)$. Let $\Psi_{z}\in{\mathcal{E}}_{-i}$ denote the normalized eigenvector corresponding to $\lambda_{z}$, and let $\Phi_{z}\in L^{2}_{K,(1,-1)}$ denote a normalized eigenvector corresponding to $\mu_{(1,-1)}(z)$, so that
(4.55) $\Psi_{z}=\psi_{1}+{\mathcal{O}}(|z|),\quad\text{ and }\quad\Phi_{z}=\phi+{\mathcal{O}}(|z|),$
where $\phi$ is defined by (3.19) with $\omega=(\zeta_{3},-1)$.
Now assume for contradiction that $\mu_{(1,-1)}(z)$ has multiplicity strictly less than 3 for $z\not=0$. Then $T\Psi_{z}$ and $T\Phi_{z}$ are also eigenvectors corresponding to $\lambda(z)$ and $\mu_{(1,-1)}$, respectively, and so both of these eigenvalues must have multiplicity at least 2. Since we are assuming that the multiplicity of $\mu_{(1,-1)}$ is strictly less than 3, we deduce that these eigenvalues must in fact be equal, and their multiplicity is exactly 2.
As a result, for all $z\in{\mathbb{R}}$, nonzero and sufficiently small, $\operatorname{span}(\Psi_{z},T\Psi_{z})=\operatorname{span}(\Phi_{z},T\Phi_{z})$. Therefore we can express $\Phi_{z}$ with respect to $\Psi_{z}$ and $T\Psi_{z}$ as
(4.56) $\Phi_{z}=\langle\Phi_{z},\Psi_{z}\rangle\Psi_{z}+\langle\Phi_{z},T\Psi_{z}\rangle T\Psi_{z},$
where we have used the fact that $T\Psi_{z}\in{\mathcal{E}}_{i}$, and is therefore orthogonal to $\Psi_{z}$. Taking the limit of both sides of (4.56) as $z\rightarrow 0$, we obtain
(4.57) $\phi=\langle\phi,\psi_{1}\rangle\psi_{1}+\langle\phi,T\psi_{1}\rangle T\psi_{1}.$
This is not possible: by (3.19), the left-hand side depends on $x_{3}$, while the right-hand side depends only on $x_{1},x_{2}$. We conclude that $H_{z}$ has a double and a triple $L^{2}_{K}$-eigenvalue for all $z\in{\mathbb{R}}$ away from a discrete set, and the corresponding eigenprojector is analytic on some open set $U$.
(3) Computation of the characteristic polynomial: Fix some $z\in U\cap{\mathbb{R}}$, and let $\phi_{1}\in L^{2}_{K,(\zeta_{3},1)}$, $\phi_{2}\in L^{2}_{K,(\overline{\zeta_{3}},1)}$ be normalized eigenvectors for the eigenvalue $\mu_{(\zeta_{3},1)}(z)$ of $H_{z}$. The entries of $M(z,\kappa)$ with respect to this basis are given by $-2i\kappa\cdot\langle\phi_{j},\nabla\phi_{\ell}\rangle$, and by Lemma 3.5, the entries of $M(z,\kappa)$ are entirely determined by $\langle\phi_{1},\nabla\phi_{2}\rangle$. However, this same lemma also tells us that $\langle\phi_{1},\nabla\phi_{2}\rangle$ is an eigenvector of $f$ with eigenvalue 1, and therefore must be the zero vector. It follows that $M(z,\kappa)=0$ for all $z$ and $\kappa$, and thus we conclude that $(K,\mu_{(\zeta,1)}(z)$ is a 2-fold quadratic point.
Now, let $\phi_{1}\in L^{2}_{K,(1,-1)}$, $\phi_{2}\in L^{2}_{K,(\zeta_{3},-1)}$, and $\phi_{3}\in L^{2}_{K,(\overline{\zeta_{3}},-1)}$ be normalized eigenvectors for the eigenvalue $\mu_{(1,-1)}(z)$ of $H_{z}$. Then the same argument implies that, for $j,\ell\in\{1,2,3\}$, $j\not\ell$̸, $\langle\phi_{1},\nabla\phi_{2}\rangle$ is again an eigenvector of $f$ with eigenvalue 1 and therefore must be the zero vector. We thus conclude that $(K,\mu_{(1,-1)}(z)$ is a 3-fold quadratic point.
4.6. Proof of Theorem 2 for the Face-Centered Cubic
Let $\Lambda=\Lambda^{FC}$, and let $(k_{1},k_{2},k_{3})$, ${\mathcal{B}}$, $K$, $m$, $G_{0}$, and ${\mathbb{U}}$ be the corresponding objects listed in the final column of Table 4. We also let $V$ be a $\Lambda$-invariant potential and let $H_{z}=-\Delta+zV$. Lastly we will again need the group element $s_{0}$ defined in (4.4).
(1) Upper bound on multiplicity: We start by computing relations among the Fourier coefficients $V_{m(j)}$ for $j\in{\mathbb{J}}$. In particular, we compute that
(4.58) $s_{0}^{-1}K$ $=K-k_{1},$
(4.59) $s_{0}^{-2}K$ $=K-k_{1}-k_{2}+k_{3},$
(4.60) $s_{0}^{-3}K$ $=K-k_{2}.$
Again using the fact that $V$ is even and invariant under $r$, it follows that for $\omega\in{\mathbb{U}}$,
(4.61) $\mu^{\prime}_{\omega}(0)=V_{0,0,0}+\omega V_{1,0,0}+\omega^{2}V_{1,1,-1}+\omega^{3}V_{0,1,0}=V_{0,0,0}+(\omega+\omega^{3})V_{1,0,0}+\omega^{2}V_{1,1,-1}.$
We then plug $\omega$ into this formula for each $\omega\in{\mathbb{U}}$ to obtain:
(4.62) $\mu^{\prime}_{1}(0)$ $=V_{0,0,0}+2V_{1,0,0}+V_{1,1,-1},$
(4.63) $\mu^{\prime}_{i}(0)=\mu^{\prime}_{-i}(0)$ $=V_{0,0,0}-V_{1,1,-1},$
(4.64) $\mu^{\prime}_{-1}(0)$ $=V_{0,0,0}-2V_{1,0,0}+V_{1,1,-1}.$
The set where the right-hand sides of the above three equations fail to be distinct is a union of three hyperplanes. Therefore we conclude that for $V$ in a set of codimension 1, the eigenvalue $||K||^{2}$ of $-\Delta$ splits into at least two simple eigenvalues and an eigenvalue of multiplicity at most two.
(2) Lower bound on multiplicity: Again let $T$ be the conjugate-parity operator and let $\phi_{1}$ be normalized eigenvector of $H_{z}$ in $L^{2}_{K,i}$. For $z\in{\mathbb{R}}$, $\phi_{2}=T\phi_{1}\in L^{2}_{K}$ is also an eigenvector of $H_{z}$ for the same eigenvalue, and
(4.65) $(s_{0})_{*}\phi_{2}(x)=\overline{\phi_{1}(-s_{0}^{\top}x)}=\overline{i\phi_{1}(-x)}=-i\phi_{1}(x),$
which implies $\phi_{2}\in L^{2}_{K,-i}$. Therefore $H_{z}$ has a double $L^{2}_{K}$-eigenvalue for all $z\in{\mathbb{R}}$ away from a discrete set, and the corresponding eigenprojector is analytic on some open set $U$.
(3) Computation of the characteristic polynomial: Fix some $z\in U\cap{\mathbb{R}}$, and let $\phi_{1}\in L^{2}_{K,i}$, $\phi_{2}\in L^{2}_{K,-i}$ be normalized eigenvectors for the eigenvalue $\mu_{i}(z)$ of $H_{z}$. The entries of $M(z,\kappa)$ with respect to this basis are given by $-2i\kappa\cdot\langle\phi_{j},\nabla\phi_{\ell}\rangle$, and by Lemma 3.5, the entries of $M(z,\kappa)$ are entirely determined by $\langle\phi_{1},\nabla\phi_{2}\rangle$. This same lemma also tells us that $\langle\phi_{1},\nabla\phi_{2}\rangle$ is an eigenvector of $s_{0}$ with eigenvalue -1, and is therefore of the form $\alpha e_{3}$ for some $\alpha\in{\mathbb{C}}$.
Thus, with respect to the basis $\phi_{1},\phi_{2}$, $M(z,\kappa)$ is given by
(4.66) $M(\kappa)=-2i\begin{pmatrix}0&\alpha\kappa_{3}\\ -\overline{\alpha}\kappa_{3}&0\\ \end{pmatrix},$
and its characteristic polynomial is $\mu^{2}-4|\alpha|^{2}\kappa_{3}^{2}$. By Lemma 3.4, the coefficient $|\alpha|^{2}$ is analytic in $z$, and therefore will be nonzero away from a discrete set if it is nonzero for $z$ sufficiently small. However, by Lemma 3.1 we can assume that, for $z$ sufficiently small, $\phi_{1},\phi_{2}$ are given by by:
(4.67) $\phi_{1}(x)$ $=e^{iK\cdot x}+ie^{iS_{0}^{3}K\cdot x}-e^{is_{0}^{2}K\cdot x}-ie^{is_{0}K\cdot x}+{\mathcal{O}}(|z|),$
(4.68) $\phi_{2}(x)$ $=e^{iK\cdot x}-ie^{is_{0}^{3}K\cdot x}-e^{is_{0}^{2}K\cdot x}+ie^{is_{0}K\cdot x}+{\mathcal{O}}(|z|).$
It follows that, for $z$ small,
(4.69) $\alpha=e_{3}\cdot\langle\phi_{1},\nabla\phi_{2}\rangle$ $=\frac{i}{4}e_{3}\cdot\left(K-s_{0}^{3}K+s_{0}^{2}K-s_{0}K\right)+{\mathcal{O}}(|z|^{2})=\pi i+{\mathcal{O}}(|z|^{2}).$
Therefore $|\alpha|^{2}$ is nonzero for $z$ sufficiently small, and thus remains nonzero on $U\cap{\mathbb{R}}$ away from a discrete set, and consequently is nonzero on ${\mathbb{R}}$ away from a closed and countable set. By Definition 1, we conclude that $(K,\mu_{i}(z))$ is a valley point.
Appendix A Appendix: Spectral Theory of the Laplacian on $L^{2}_{K}$
In this appendix, we compute the spectrum of of $-\Delta$ on $L^{2}_{K}$ and show that the cardinality of the set $[K]$, defined by (3.13), is equal to the multiplicity of $||K||^{2}$ as an eigenvalue. We then use this to compute some bounds on the multiplicity of $||K||^{2}$ when $K\in{\mathcal{B}}$, and lastly show that if $K$ is a vertex of ${\mathcal{B}}$, then $[K]$ is a subset of the vertices of ${\mathcal{B}}$.
Fix some lattice $\Lambda$ with basis $v_{1},\ldots,v_{n}$ and reciprocal basis $k_{1},\ldots,k_{n}$, and fix some $K\in{\mathbb{R}}^{n}$. Just as we did following (3.10), we also let $mk=m_{1}k_{1}+\cdots m_{n}k_{n}$ for $m\in{\mathbb{Z}}^{n}$. We then claim that $\phi_{m}(x)=e^{i(K+mk)\cdot x}$ for $m\in{\mathbb{Z}}^{n}$ is an orthonormal basis of eigenvectors for $-\Delta$ on $L^{2}_{K}$. Indeed, note that
(A.1) $-\Delta\phi_{m}(x)=||K+mk||^{2}\phi_{m}(x),$
and $(\phi_{m})_{m\in{\mathbb{Z}}^{n}}$ is the image of the orthonormal basis $(\oplus_{j=1}^{n}e^{2\pi im_{j}x_{j}})_{m\in{\mathbb{Z}}^{n}}$ of $L^{2}[0,1]^{\otimes n}$ under the unitary map which first sends $\oplus_{j=1}^{n}e^{2\pi im_{j}x_{j}}$ to $e^{imk\cdot x}\in L^{2}_{0}$, and then $e^{imk\cdot x}$ to $\phi_{m}$ via multiplication by $e^{iK\cdot x}$. Consequently,
(A.2) $\sigma(-\Delta)=\{||K+mk||^{2}\;:\;m\in{\mathbb{Z}}^{n}\}.$
and the multiplicity of an eigenvalue $\mu_{m}:=||K+mk||^{2}$ is given by
(A.3) $m_{-\Delta}(\mu_{m})=\left|\left\{k^{\prime}\in K+\Lambda^{*}\;:\;||k^{\prime}||^{2}=||K+mk||^{2}\right\}\right|.$
In particular, this proves (3.13); namely that the the cardinality of the set $[K]$ is equal to the multiplicity of $\mu_{0}=||K||^{2}$.
From here, recall that the Floquet–Bloch problem (3.4) is periodic with respect to the dual lattice $\Lambda^{*}$, and so we focus our analysis on $K\in{\mathcal{B}}$. For such $K$, the minimal eigenvalue of $-\Delta$ on $L^{2}_{K}$ is then given by $\mu_{0}$, since by definition of the Brillouin zone ${\mathcal{B}}$,
(A.4) $||K||^{2}\leq||K-k^{\prime}||^{2},\quad\forall k^{\prime}\in\Lambda^{*}.$
This also implies that if $K$ is in the interior of ${\mathcal{B}}$, then the inequality (A.4) is in fact strict, and so the eigenvalue $\mu_{0}$ is simple. Conversely, we expect $K\in{\partial}{\mathcal{B}}$, and in particular the vertices of ${\mathcal{B}}$, to correspond to eigenvalues of high multiplicity, as the following proposition demonstrates.
Proposition A.1.
Let $K\in{\mathcal{B}}$, let $\mu_{0}=||K||^{2}$, and let $m$ be the number of (hyper)faces of ${\mathcal{B}}$ which contain $K$, where $m$ is possibly zero. Then there exist vectors $K_{1},\ldots,K_{m}\in\Lambda^{*}$ such that $K-K_{j}$ also lies on $m$ (hyper)faces of ${\mathcal{B}}$ and $||K-K_{j}||^{2}=||K||^{2}$ for $j=1,\ldots,m$, so that $m_{-\Delta}(\mu_{0})\geq m+1$. Furthermore, $m_{-\Delta}(\mu_{0})=1$ if and only if $m=0$ and $m_{-\Delta}(\mu_{0})=2$ if and only if $m=1$.
Proof.
Let $K\in{\mathcal{B}}$ such that $K$ lies on $m$ (hyper)faces of ${\mathcal{B}}$ for some non-negative integer $m$. Then if $m>0$, there exist vectors $K_{1},\ldots,K_{m}\in\Lambda^{*}$ such that $K$ lies on the (hyper)planes defined by $x\cdot K_{j}=\tfrac{1}{2}||K_{j}||^{2}$ for $j=1,\ldots,m$, where each of these (hyper)planes intersected with ${\mathcal{B}}$ is precisely one of the $m$ (hyper)faces containing $K$. Then for $j=1,\ldots,m$
(A.5) $||K-K_{j}||^{2}$ $=||K||^{2}-2K\cdot K_{j}+||K_{j}||^{2}$
(A.6) $=||K||^{2}-||K_{j}||^{2}+||K_{j}||^{2}$
(A.7) $=||K||^{2}.$
Therefore it follows from (A.3) that $\mu_{0}$ has multiplicity of at least $m+1$.
To prove that $K-K_{j}$ lies on $m$ (hyper)faces of ${\mathcal{B}}$, observe that since $K\in{\mathcal{B}}$, it follows that,
(A.8) $||K-K_{j}||^{2}=||K||^{2}\leq||(K-K_{j})-K||^{2},\quad\forall K\in\Lambda^{*},$
which implies $K-K_{j}\in{\mathcal{B}}$. Furthermore, we have that
(A.9) $(K-K_{j})\cdot(-K_{j})=-K\cdot K_{j}+||K_{j}||^{2}=-\frac{1}{2}||K_{j}||^{2}-||K_{j}||^{2}=\frac{1}{2}||K_{j}||^{2},$
and for all $\ell\not=j$,
(A.10) $(K-K_{j})\cdot(K_{\ell}-K_{j})$ $=K\cdot K_{\ell}-K\cdot K_{j}-K_{j}\cdot K_{\ell}+||K_{j}||^{2}$
(A.11) $=\frac{1}{2}||K_{\ell}||^{2}-\frac{1}{2}||K_{j}||^{2}-K_{j}\cdot K_{\ell}+||K_{j}||^{2}$
(A.12) $=\frac{1}{2}\left(||K_{\ell}||^{2}-2K_{j}\cdot K_{\ell}+||K_{j}||^{2}\right)$
(A.13) $=\frac{1}{2}||K_{\ell}-K_{j}||^{2}.$
Therefore $K-K_{j}$ lies on the $m$ (hyper)planes defined by $x\cdot(-K_{j})=\tfrac{1}{2}||K_{j}||^{2}$ and $x\cdot(K_{j}-K_{\ell})=\tfrac{1}{2}||K_{j}-K_{\ell}||^{2}$ for $\ell\not=j$.
We now seek to show that each of these (hyper)planes defines a (hyper)face of ${\mathcal{B}}$. To start, for $j=0,\ldots,m$ and $\ell=1,\ldots,m$, let
(A.14) $P_{j\ell}=\begin{cases}\{x\in{\mathbb{R}}^{n}\;:\;x\cdot K_{\ell}=\tfrac{1}{2}||K_{\ell}||^{2}\}&j=0\\ \{x\in{\mathbb{R}}^{n}\;:\;-x\cdot K_{j}=\tfrac{1}{2}||K_{j}||^{2}\}&j\not=0,\ell=j\\ \{x\in{\mathbb{R}}^{n}\;:\;x\cdot(K_{\ell}-K_{j})=\tfrac{1}{2}||K_{j}-K_{\ell}||^{2}\}&j\not=0,\ell\not=j,\end{cases}$
and suppose $k^{\prime}\in P_{0\ell}\cap{\mathcal{B}}$ for some $\ell$. Then the same computations as in (A.5)-(A.8), but with $K$ replaced with $k^{\prime}$, imply that $k^{\prime}-K_{j}\in{\mathcal{B}}$ for $j=1,\ldots,m$. Similarly, (A.9) with $K$ replaced with $k^{\prime}$ implies $k^{\prime}-K_{j}\in P_{jj}$, and (A.10) with $K$ replaced with $k$ implies $k^{\prime}-K_{j}\in P_{j\ell}$ for $j\not=\ell,0$. As a result, for $j=1,\ldots,m$,
(A.15) $(P_{0\ell}\cap{\mathcal{B}})-K_{j}=P_{j\ell}\cap{\mathcal{B}}.$
By construction though, $P_{0\ell}\cap{\mathcal{B}}$ is a (hyper)face of ${\mathcal{B}}$, and since $P_{j\ell}\cap{\mathcal{B}}$ is an isometric set and must be contained in the boundary of ${\mathcal{B}}$, it follows that $P_{j\ell}\cap{\mathcal{B}}$ is in fact a (hyper)face of ${\mathcal{B}}$ as well.
For the second part of the proposition statement, observe that it suffices to prove that $m_{-\Delta}(\mu_{0})\leq m+1$ when $m=0,1$. However, if $m=0$ then this means that $K$ lies on zero (hyper)faces, and therefore must be in the interior of ${\mathcal{B}}$. We have already seen that in this case the eigenvalue $\mu_{0}=||K_{0}||^{2}$ is simple, and thus $m_{-\Delta}(\mu_{0})=1$, as desired. Now assume that $m=1$, so that $K$ lies on a single (hyper)face of ${\mathcal{B}}$, and let $m^{\prime}=m_{-\Delta}(\mu_{0})$. Then by again using (A.3), we deduce that there exist vectors $K_{1},\ldots,K_{m^{\prime}}\in\Lambda^{*}$ such that $||K-K_{j}||^{2}=||K||^{2}$ for $j=1,\ldots,m$. As a result, (A.10) implies that $K\cdot K_{j}=\tfrac{1}{2}||K_{j}||^{2}$ for $j=1,\ldots,m^{\prime}$, and so $K$ lies on the $m^{\prime}$ distinct (hyper)planes defined by $x\cdot K_{j}=\tfrac{1}{2}||K_{j}||^{2}$ for $j=1,\ldots,m^{\prime}$. However, since $K$ lies on a single (hyper)face of ${\mathcal{B}}$, this implies $K$ lies on exactly one of these (hyper)planes. Therefore $m^{\prime}=m_{-\Delta}(\mu_{0})=1$, as claimed. ∎
Proposition A.2.
Let $V({\mathcal{B}})$ denote the vertices of ${\mathcal{B}}$, let $K\in V({\mathcal{B}})$, and let $\mu_{0}=||K||^{2}$. Then
(A.16) $[K]=V({\mathcal{B}})\cap(K_{0}+\Lambda^{*}).$
Proof.
Let $K\in V({\mathcal{B}})$ and let $k^{\prime}\in\Lambda^{*}$; it then suffices to prove that $||K-k^{\prime}||^{2}=||K||^{2}$ if and only if $K-k^{\prime}\in V({\mathcal{B}})$. However, (A.5) implies that $||K-k^{\prime}||^{2}=||K||^{2}$ if and only if $K\cdot k^{\prime}=\tfrac{1}{2}||k^{\prime}||^{2}$, and so by the proof of Proposition A.1, $K-k^{\prime}$ lies on the same number of (hyper)faces of ${\mathcal{B}}$ as $K$ does. Together with the fact that $K\in V({\mathcal{B}})$, this implies $K-k^{\prime}\in V({\mathcal{B}})$ as well. ∎
[…]
[heading=bibintoc]